A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.08

Title word cross-reference

(0, 1) [24]. (1, 1) [377]. (G, 2) [328]. (N, p, q) [347]. (p^{2} − 1) [96]. (T, M, S) [652]. (V^{∗}) [467]. (X, L) [629]. ∗ [23]. 0 [534, 309, 510, 581]. {0, +1, −1} [551]. 1 [534, 565]. 17 [92]. 1: 2 [11]. 2 [75, 387, 43, 430, 11, 466, 518]. 2m [88]. 2 × 2 [331]. 3 [76, 153, 144, 287]. 32 [341]. 3x + 1 [494]. 4 [552, 296]. 5 [514]. 7/10 [29]. [−1, 1] [632, 660]. A_{∞}^{+} [285]. A_{p}^{0} [368]. A^{n} [502]. A^{n−1} [502]. A_{N} [320]. A_{p}(G) [501]. A_{ϕ} [576]. B(Z/p)_{+}^{n} [143]. ¯∂ [568]. βN [334]. (U(p, q), U(r, s)) [664]. C(α) [651]. C(X) [319, 554]. C[−1, 1] [194]. C^{∗} [37, 667, 156, 485, 562, 159, 554, 336, 441, 338, 442, 148, 10]. C^{1} [433]. C^{2} [607]. C^{∗} [470, 356, 587, 349]. C^{∞} [477]. C_{p} × C_{p} [539]. M [535]. DF [625]. E_{0} [405]. epsilon [423]. F [625]. F_{4} [653]. FC [380]. G [501]. G(r, p, n) [579]. G^{∘} [385]. G_{2} [500]. G_{2} × PU_{3}(D) [643]. GL(3) [210, 284]. GL(n) [306]. GL_{2} [453]. Gsb2(q) [469]. G ⊆ SU_{2}(C) [689]. H [328, 580]. H^{2} [72, 163]. H^{∞} [185, 644]. H_{1} [433]. II_{1} [68]. JW [406]. K [267, 483, 688]. K[![x, y]!] [407]. K_{2} [234]. K_{3} [174]. K_{X}L [629]. L [685, 28, 541, 671]. l^{1} [467]. l^{2} [9, 229]. l_{p}^{n} [321]. L^{p} [507, 660, 540, 530]. L_{1} [173]. L_{2}(R^{d}) [422]. L_{p} [630]. LIMsimTLIM [337]. n [591, 14]. n ≥ p [177]. o [187]. oldQ(N^{2} + 4) [420]. p [385, 387, 59, 631, 95, 249, 489, 643, 665]. P (n) [488]. p, q [393]. P^{2} [147]. p^{n} [400, 19]. p_{−k}(n) [92]. PGL(3) [59]. pi [498, 84]. p ≠ 2 [106]. pq [372]. q [436, 598, 535, 529, 16, 390, 592, 189]. Q_{p} [594]. R [615]. R^{3} [396]. R^{d} [512]. R^{n} [51, 66, 410]. RG [19]. S_{R}^{−p} [33]. S_{k}(n, p) [177]. S_{N} [320]. scrA [348]. scrK [348]. SK_{2} [174]. SL(3, R) [552]. SL(n, R)/ SO(n) [522]. SO_{2n+1} × SO_{2m} [344]. SO_{n} [197]. SO_{n−1} [197]. SU(2) [390]. T_{R}^{−p} [33]. Θ [590]. θ(z, q) [254]. U(3) [398]. U(n, n) [645]. U_{t}(rmsl(2)) [542]. VN(G) [559]. W [581]. W^{∗} [510]. y′′ = f(t, y, y′) [117, 117]. Z_{2} [483]. Z_{p} [631, 605]. |L(1, χ)| [612]. |x|^{α} [556].

-Adic [643, 385, 59]. -algebra [562]. -algebras [23, 667, 470, 587, 406, 485, 37, 554, 336, 441, 338, 442, 148, 10]. -analogues [393]. -Bessel [189]. -bimodules [356]. -Carleson [535]. -centre [591]. -class [631]. -complemented [565]. -convex [349]. -convexity [156]. -dimensional [552, 76, 309, 287, 14, 96]. -extensions [43, 631]. -families [423]. -Functions [685, 671, 541]. -group [380]. -Groups [665, 430, 489]. -harmonic [535]. -harmonics [580]. -Hermite [436]. -invariant [576]. -Kostka [598]. -Laguerre [16]. -lattice [19]. -lattices [24]. -Legendre [390]. -local [615]. -manifolds [514, 153, 144]. -means [347]. -minimal [187]. -Nets [652]. -Neumann [568]. -norm [9, 229]. -partial [498]. -power [625]. -rank [518]. -ranks [296]. -regular [95, 249]. -representations [590]. -second [477]. -semigroups [405]. -separable [84]. -series [529, 592, 28]. -solvable [95, 249]. -space [368]. -spaces [630]. -structures [328]. -subalgebras [159]. -Sylow-subgroup [75]. -tactics [267]. -theoretic [483]. -Theory [688]. -tic [400]. -unconditional [565].

A. [150]. a.s [286]. abelian [354, 62, 481, 80, 376, 681, 663]. absolute [112]. Abstract [282, 360]. across [411]. Action [512, 473, 283]. actions [483, 452, 69, 252, 605]. Acyclic [447]. Acyclicity [5]. Addendum [60, 150]. Addition [379, 390, 460, 189]. Additive [316, 101, 26, 340]. Adic [643, 385, 59]. adjoint [571, 511]. adjunction [536]. adjunction-theoretic [536]. advances [420]. AF [483, 79, 389, 69]. AF-algebras [79]. Affine [689]. Algebra [684, 673, 413, 448, 360, 269, 463, 559, 562, 262, 211, 532, 489]. Algebraic [354, 658, 359, 330, 187, 326, 547, 615, 226, 633]. Algebras [657, 667, 659, 672, 665, 457, 435, 23, 582, 470, 255, 618, 279, 177, 585, 15, 587, 483, 343, 257, 207, 332, 269, 184, 589, 270, 579, 79, 389, 406, 485, 373, 37, 554, 336, 41, 639, 510, 441, 272, 311, 160, 338, 312, 442, 374, 21, 237, 566, 188, 96, 148, 10, 581, 487]. Algèbres [639, 546]. algébriques [226]. algorithm [407, 2]. almost [595, 493, 227, 609, 444]. along [150]. Alternating [106]. alternative [519]. amalgamated [336]. amalgams [22]. ambiguity [82]. ambiguous [620]. amenability [362]. amenable [622, 289]. ample [150]. ampleness [492]. analogues [254, 395, 393]. Analysis [666, 185, 179, 181, 277]. Analytic [392, 428, 37, 272, 311, 644, 180, 401, 369, 449, 155, 264, 564, 456]. anisotropic [544]. appear [563]. applicable [266]. Application [597, 671, 179, 227, 116, 361, 11, 125]. Applications [650, 76, 359, 330, 22, 168, 305, 198, 69, 612, 454, 397, 642, 14, 97]. approach [162]. approximants [556]. approximate [36, 312]. approximating [597]. Approximation [194, 195, 553, 459, 183, 258, 508, 118, 324, 530, 214, 642]. approximations [248]. argument [181]. arising [367, 587, 485]. arithmetic [601]. Arrays [652, 519]. Artamonov [305]. Artin [149]. Aspects [657]. asphericity [232]. Associated [133, 665, 470, 208, 35, 173, 408, 580].

Association [652]. Asymptotic [647, 572, 592, 265, 490, 366, 642, 123]. Asymptotics [361, 105]. ATM [361]. atomic [598, 220]. Atomicity [20]. attaining [383]. attracting [367]. attractive [520]. auteurs [570, 635, 696]. Author [570, 635, 696]. Automorphic [677, 671, 316]. Automorphism [102, 279, 68, 590]. Automorphisms [15, 50, 510, 470, 241, 292, 387, 21]. Averages [660, 239, 384, 410, 276]. avoid [142]. axes [44]. axiom [109]. axioms [67].

Bailey [32]. baker [251]. balanced [100]. Banach [267, 381, 638, 661, 476, 54, 248, 160, 289, 291, 188, 191]. Bargmann [670]. based [247]. Bases [688, 257, 565, 422, 298]. basic [236]. BC [523]. be [56]. Beauregard [217]. behavior [118, 265, 297]. behaviour [416, 630, 642, 123]. Berg [69]. Bergman [445, 482]. Bernstein [91, 416, 250]. Bessel [61, 49, 136, 189, 98]. Best [214, 183, 118, 556]. between [429, 271, 581]. Bi [668]. Bi-Period [668]. Bianchi [70]. bibasic [1]. bicubic [25]. Bicyclic [25, 518]. bifurcations [11, 521]. big [390]. billiard [597]. bimodules [356]. binômes [218]. binary [379, 282]. binomials [218]. biquadratic [620, 518]. Birelative [234]. Birkhoff [212]. bivariant [164]. Blaschke [656]. Bloch [316]. block [100, 200, 96]. blocks [414]. Bochner [467]. bodies [3]. body [129]. Boltzmann [115, 208]. Boolean [73, 373]. Borel [204]. bound [629, 363]. boundaries [152, 333]. boundary [514, 129, 158, 388, 440]. Bounded [459, 45]. Boundedness [280, 544, 114, 362]. Bounds [699, 119, 140, 599, 612, 226]. Bourgain [184]. box [63]. braid [223]. branched [144, 150]. branching [506, 295, 116]. Brauer [414, 497, 395, 13]. Bridge [357]. Brown [335]. Brownian [666, 462, 123]. Bruhat [193]. Brunn [655]. bump [248]. bundle [253]. bundles [492, 39, 160]. Butler [65].

C [150]. calculus [476, 266]. can [56]. cancellation [31]. canonical [309]. Carathéodory [244, 117]. Carleson [426, 535, 209]. Carnot [244]. Carpets [666, 682]. carrés [639]. Cartan [581]. cascades [103]. Case [643, 367, 514, 609]. cases [258]. categories [12]. category [107, 615]. Cauchy [603, 120]. cells [397]. center [531]. Central [457, 192, 29, 172, 441]. centralizer [503]. centralizers [219]. Centre [673, 591]. CEP [104]. certain [180, 282, 52, 143, 497, 464, 347, 50, 622, 58, 420, 55, 5]. certains [622]. chain [325]. chains [381, 142]. character [403, 398, 164]. Characteristic [525, 683, 43, 510, 581]. characteristics [469]. characterization [178, 368, 93, 555]. characterizations [360, 321, 519]. characterize [153]. Characters [641, 579, 498, 84]. charge [29]. Chébli [617]. Chern [164]. Chevalley [469]. Chinburg [307]. chiral [399]. chromatic [293]. Circle [545, 30]. Circulant [296]. circular [382]. Class [186, 139, 619, 307, 296, 631, 60, 18, 299, 137, 518, 251, 489, 490, 34, 566, 216]. class-group [566]. class-numbers [60, 18]. Classes [396, 446, 358, 359, 330, 33, 39, 95, 249, 326, 614, 273]. Classical [528, 512, 436, 135]. Classification [220, 548, 76, 483, 402, 82]. classifications [43]. Classifying [514]. Cliques [414]. clone [503]. clones [345]. close [551]. Closed [413, 502]. closure [124]. Co [328]. Co- [328]. Coactions [667]. cobordism [317, 252]. Cocharacters [331]. cocycles [242]. Codimension [11, 342, 536]. codimensions [331]. coefficients [551]. coercive [408]. Cohomology [674, 288, 76, 211, 441]. coinvariants [261]. Collapsing [244]. Collision [111]. Coloring [142]. combinaisons [226]. combinations [226]. Combinatorial [685, 213, 237]. Commutative [160, 166, 639]. commutatives [639]. Commutativity [255]. commutators [102]. Commuting [41, 275]. Compact [670, 444, 367, 452, 559, 45, 362, 501, 235, 55]. compacta [298]. compactly [63]. Compactness [208, 501]. Comparison [394]. compartmental [125]. Competing [654]. Complemented [467, 565]. complements [232]. completable [315]. complete [520, 21, 147]. completely [178, 42]. completeness [187]. completions [477]. Complex [674, 135, 342, 193, 8, 579, 429, 211, 642, 411, 616, 253]. complexes [8, 107]. complexity [187, 85]. components [548, 99]. Composite [648]. composition [45]. Compound [625]. concerning [420]. Condition [281, 285]. Conditional [238, 265]. conditions [471, 472, 131, 299, 603, 325, 544]. conductor [147]. cones [191]. Configurations [693]. congruence [73]. congruences [292, 92]. Conjecture [636, 469, 382, 212, 550]. conjectures [420, 203]. Conjugacy [359, 330, 326, 614, 494]. conjugate [7, 604]. connected [385, 38, 487]. Connes [68]. Consistency [418]. constant [181, 561]. constants [524]. Constrained [508]. constraints [214]. Construction [434]. Constructions [13, 154, 587]. constructive [497]. Contact [684, 260, 477]. Contactomorphisms [675]. containing [345, 119]. continua [357]. Continuation [411]. Continuations [161, 564]. continued [412, 186, 274]. continuity [112, 9, 229]. Continuous [181, 401, 76, 133]. Continuum [246, 109]. Continuum-wise [246]. Contractions [540]. contractive [312]. contrasting [313]. contre [583]. Contributions [243]. controllability [277]. Convergence [648, 112, 114, 471, 472, 116, 632, 520, 297].

Converse [660]. Convex

[675, 152, 3, 89, 278, 198, 81, 349, 456]. Convexity [281, 156, 333, 347, 190]. Convexly [648]. Convolution [377, 413, 63, 299, 364, 393]. coordinates [123]. copies [467]. corona [10]. Correction [359, 699, 249, 311]. correspondence [623]. correspondences [403, 84]. Corrigenda [61]. Corrigendum [697, 698]. Corson [298]. coset [463]. cotype [466]. countably [289]. counterexample [583, 530]. counterexamples [375]. Counting [699, 679, 599]. Covering [267, 106]. coverings [457, 134, 144, 376, 365, 150]. Covers [681, 186]. Coxeter [695, 689]. Criteria [281, 492, 478, 276]. criterion [152, 116]. critical [402, 577, 199, 521]. crochet [302]. cross [619]. cross-sections [619]. crossed [485, 441, 273]. crossing [223]. Cubic [254, 1, 236]. Curvature [247, 181, 437, 52, 561]. curve [492]. Curves [683, 264, 212, 122]. cusps [396]. CW [107]. CW-complexes [107]. cycles [354, 525]. cyclic [586, 124, 150].

D. [150]. Dade [469]. dans [396]. d’aronde [396]. dealing [287]. Decaying [88]. deciding [407]. decomposable [339]. Decomposition [167, 618, 431, 672, 598, 197]. decompositions [563]. Dedekind [187, 646, 698]. defect [307]. defined [515, 215]. defining [469, 464]. definite [74, 371, 479, 443, 175]. deformations [152]. Degenerate [258]. degenerations [491]. Degree [324, 155, 40]. degrees [577]. delay [394]. Dense [279, 270]. Densities [373]. Density [538, 456, 53, 47]. dependence [182]. dependent [138]. derivation [50, 172]. Derivations [582, 241, 219, 121]. derivative [129]. derivatives [266]. designs [100]. Determinantal [517, 34]. determined [558]. deux [120]. d’homotopie [302, 396]. diagrams [40, 107]. Diameters [3]. Dickson [300]. difference [446]. different [358, 43]. differentiability [278]. differentiable [248, 191]. Differential [180, 515, 181, 493, 304, 171, 478, 224, 313, 276, 125]. differentiation [27]. différentiel [583]. Diffusion [314, 352, 468]. diffusions [367, 295]. dihedral [174, 365]. Dilatations [656]. dilation [391]. Dilations [540, 41]. dimension [629, 288, 546]. Dimensional [662, 552, 76, 542, 93, 309, 287, 490, 225, 313, 14, 96]. Dimensions [692, 699, 682, 129, 599]. Diophantine [105]. direct [325, 338]. directed [230]. Dirichlet [228, 375, 291]. disc [585]. Disciplined [503]. disconnected [445]. Discrete [667, 564, 444]. Disjointly [26]. disproof [420]. Distance [409, 290, 610]. distinct [627]. Distribution [545, 315, 127, 63, 438, 361, 286]. distribution-valued [127]. distributions [36, 113, 17, 66, 14]. distributive [292]. distributivity [128]. divergence [596]. divergent [85]. divisibility [230]. division [271]. divisor [150]. divisors [492]. Dixmier [273]. domain [568]. Domains [697, 445, 516, 65, 388, 397, 607]. Douady [273]. double [463, 504]. Doubling [154, 103]. Douglas [184]. dovetails [396]. Dual [664, 643, 397, 455]. duality [480, 484, 298]. Dyer [269]. dynamical [273]. Dynamics [512, 531]. Dynkin [689].

each [48]. Effective [560]. effectives [560]. Eichler [605]. Eigenpolytopes [610]. Eigenvalue [557, 139]. Eigenvalues [52]. Einstein [253]. Eisenstein [645]. Elementary [381, 575]. Elements [688, 619, 227, 282, 628, 417, 56, 121]. elliptic [88, 492, 459, 264, 199, 108]. Embeddings [662, 51, 356, 69, 213, 502, 616]. Endoscopic [685]. endoscopy [417]. Ends [697, 516]. energy [199]. entières [218]. entire [368, 64, 474]. entropy [151, 238]. Enumerator [647]. envelopes [374]. enveloping [589]. epi [266]. epi-derivatives [266]. épimorphismes [245]. epimorphisms [245]. equation [171, 560, 287, 547, 175, 560]. Equations [680, 36, 595, 493, 459, 515, 17, 8, 304, 115, 208, 78, 198, 478, 263, 224, 108, 313, 276, 125, 87, 277, 544, 8, 115]. equidecomposability [31, 353]. equivalence [567, 302]. equivalences [6, 302]. Equivariant [103, 590, 283]. Erdos [471, 472, 47]. Ergodic [679, 410, 384, 404, 308, 54, 297, 351, 190]. Erratum [700]. espaces [622]. essential [481]. Essentially [275]. estimate [192]. Estimates [611, 507, 649, 522, 377, 446, 626, 129, 202, 149]. Eta [646, 698]. Euclidean [690]. Euler [204]. every [56]. everywhere [85]. Exact [329, 40, 624]. exactly [282]. examples [375]. Exceptional [658, 653]. exchangeable [182]. exclude [201]. exemple [583]. Existence [694, 117, 289, 597, 402, 117]. exit [361]. expansion [490]. expansions [379, 131, 428, 398]. expansive [470, 246]. Explicit [453, 43, 122, 567, 568]. exponent [458, 78]. Exponential [362, 424, 64, 632, 338]. exposed [169]. extension [124]. Extensions [476, 435, 481, 539, 620, 43, 631, 337, 399, 238]. Exterior [656, 671, 511]. Exton [189]. Extra [341]. Extra-special [341]. extrema [366]. Extremal [33, 165]. Extreme [678, 415, 169, 188].

facially [220]. factor [68]. factorisability [307]. Factorization [389, 562, 55]. Factors [668, 329, 205]. failure [31]. families [354, 35, 201, 423]. Family [513, 450, 256, 57, 264, 400]. Fan [183]. Fejér [173]. Fermat [218, 560]. Fiber [477]. Fibers [675]. fibrations [588]. field [59, 43, 124]. Fields [637, 75, 425, 318, 515, 262, 60, 18, 186, 518, 420, 25, 400]. Fine [526, 534]. Finite [656, 495, 345, 542, 662, 481, 582, 572, 403, 283, 104, 82, 465, 363, 199, 607, 504]. finitely [340]. First [684, 526, 534, 473, 664]. First-order [526, 534]. Fixed [387, 190, 191, 176, 103, 360, 119]. flag [252, 569, 548]. Flat [180]. flow [608, 537]. flows [334]. Fluctuations [115, 208, 115]. FN [368]. fondamental [110]. Fong [84]. Forbidden [487]. forces [129]. Form [670, 596, 380, 318, 634]. formal [616]. formality [283]. forme [634]. Forms [677, 690, 180, 23, 582, 437, 282, 517, 453, 408, 375, 291, 434]. Formula [668, 584, 1, 386, 390, 567, 32]. Formulae [647, 393]. Formulas [335, 64, 1, 236, 189]. formule [32]. Four [692, 595, 100, 282, 56, 200]. Fourier [617, 22, 66, 322, 85, 297, 327]. fractal [202]. Fractals [606]. Fraction [270, 412, 274]. Fraction-dense [270]. fractional [573, 555, 454, 603]. fractions [543, 186]. Frames [200, 422]. Fréchet [278, 638]. Fréchet-Like [638]. Fredholm [608]. Free [405, 58, 399, 481, 584, 227, 279, 318, 586, 15, 587, 609, 170, 387, 234, 479, 532, 34]. freedom [577]. French [583, 218, 8, 302, 115, 117, 245, 560, 396, 622, 120, 634, 639, 546, 226, 110, 32]. Fubini [46]. Fuchsian [227, 70, 120, 524]. fuchsiennes [120]. fully [486]. fully-ramified [486]. Function [699, 646, 698, 7, 424, 599, 332, 254, 467, 290, 548, 147]. Functional [36, 17, 71, 125]. functional-differential [125]. functionals [383, 402, 171, 188]. Functions [685, 507, 626, 648, 671, 680, 545, 513, 126, 401, 316, 493, 368, 369, 74, 167, 278, 53, 413, 558, 535, 392, 358, 449, 228, 194, 303, 574, 64, 195, 474, 386, 517, 222, 306, 541, 135, 553, 101, 324, 543, 61, 49, 136, 530, 410, 248, 578, 482, 479, 351, 642, 189, 98, 216]. functors [12]. fundamental [575, 417, 93, 110]. fusion [195].

Gâteaux [278]. gages [215]. Galois [676, 539, 620, 628, 134, 395, 430, 431, 341, 365]. game [267]. games [267]. Gateaux [248]. Gaussian [113, 127, 9, 294, 71, 202]. Gel’fand [160]. General [153, 243, 204, 87]. généralisée [560]. generalization [183, 439, 16]. Generalized [22, 393, 126, 377, 378, 450, 91, 208, 116, 171, 560, 613, 397, 251, 216, 581]. Generated [686, 9, 294, 633, 289, 125]. Generators [464]. Generic [242]. Genus [700, 514, 83, 363, 533]. Geodesic [537]. Geometric [368, 501]. Geometry [303, 67, 350]. germ [155]. Gilbert [217]. given [551, 415, 48, 201, 149]. Gorenstein [527]. Goresky [636]. Goulden [550]. governed [277]. Graded [231, 582]. Gradient [626]. Gradings [657]. graph [39, 213, 265, 201]. Graphs [681, 689, 293, 627, 230, 610, 296, 201, 376]. Greenberg [499]. Green’s [574]. Gröbner [257]. Gross [382]. Group [67, 86, 252, 148, 571, 378, 380, 611, 495, 586, 130, 473, 283, 575, 372, 307, 559, 395, 390, 37, 562, 591, 543, 363, 518, 501, 2, 21, 489, 566, 504]. Group-theoretic [67]. groupes [622, 110]. Groups [686, 667, 658, 670, 134, 695, 687, 106, 665, 355, 179, 192, 469, 193, 458, 62, 481, 619, 128, 227, 279, 76, 221, 359, 330, 586, 102, 414, 403, 588, 385, 452, 170, 579, 387, 497, 260, 231, 541, 296, 362, 499, 68, 174, 430, 42, 622, 631, 70, 234, 56, 593, 431, 350, 84, 94, 95, 249, 288, 271, 86, 489, 5, 341, 34, 72, 163, 314, 352, 468, 524, 110, 580]. growth [179, 205, 544, 487]. Gyroscopically [139].

Hahn [181, 133, 189]. half [66]. half-space [66]. Hamiltonian [80, 577, 531]. Hamiltonians [633]. handlebodies [83]. Hankel [607]. Hardy [126, 617, 426, 209, 604, 222, 509, 454, 482, 351]. Harmonic [666, 656, 179, 535, 626, 303, 158, 145, 530, 275]. harmonics [580]. Harnack [179]. Hasse [541, 634]. Hauptmoduls [427]. Hausdorff [178, 322, 682]. having [278, 147]. head [286]. heat [265, 175, 522, 202]. Hecke [579, 306, 566]. Heights [28]. Heisenberg [343, 659]. Hereditarily [661]. heredity [480]. Hermite [436, 173, 640]. Hermitian [525, 547, 253]. Herz [454]. Herz-type [454]. hierarchically [367]. higher [166, 317, 224]. higher-order [224]. Hilbert [356, 331, 384, 46, 14, 147]. Hölder [78]. holomorphic [413, 388, 429, 564]. Homeomorphic [644]. homeomorphism [5]. homeomorphisms [246]. homoclinic [11]. Homogeneous [219, 342, 515, 52, 102, 230, 622, 235, 502, 253]. homogènes [622]. homogenization [179]. Homological [480, 657]. Homology [144, 681, 376, 342, 602, 261]. homomorphism [458]. Homomorphisms [332, 292, 554, 34]. Homotopy [650, 107, 458, 76, 302, 211, 6, 44, 615, 396]. hook [302]. Hopf [435, 269, 488, 237, 673]. Horn [697, 516]. Horn-Like [697, 516]. Hughes [57]. Hulls [675]. Hurewicz [458]. Hurwitz [524]. hyperbolic [499, 263]. hypercentral [86]. hypercentre [591]. hyperelliptic [122]. hypergeometric [236, 216]. Hypergroups [641, 74, 617, 444]. hypersurfaces [13]. hypoellipticity [299]. hypothesis [109].

Ideal [644, 537, 259, 296, 185]. ideals [413, 620, 464, 47, 172, 186]. idempotents [241]. identities [36, 331, 59, 172, 312]. Identity [685]. II [624, 180, 140, 141, 534, 472, 527, 64, 292, 260, 210, 101, 651, 105, 543, 233, 326, 569, 352, 164, 98]. III [284, 137, 468, 10]. images [235]. imaginary [518]. Imbedding [433]. immanants [203]. implicit [245]. implicites [245]. Imprimitively [633]. incomplete [71]. increasing [222]. Indecomposable [661, 609, 491]. Indefinite [557, 1]. independence [63]. Index [570, 635, 696, 597, 103, 385, 223]. Induced [667, 672, 90]. Induction [498]. inductive [483]. Inequalities [660, 578, 179, 126, 329, 51, 573, 91, 222, 561, 322, 555, 351]. inequality [209, 364, 250]. infinie [546]. infinite [448, 587, 452, 394, 440, 546, 225]. infinite-dimensional [225]. Infinitely [680]. infinity [191]. Inflection [136, 61, 49, 98]. injective [269, 346, 68, 325]. inner [305, 121]. integer [138, 218, 63]. integer-valued [138]. Integers [695, 205, 539]. intégrabilité [546]. integrability [546]. integrable [64, 577]. Integral [130, 631, 112, 23, 280, 573, 495, 372, 59, 591, 247, 276, 277]. intégrales [110]. integrals [370, 549, 306, 210, 93, 284, 421, 490, 110]. Intégration [583, 240, 454, 173, 583]. interacting [367]. interactions [295]. interior [158, 119]. internal [10]. Interpolants [678]. interpolating [368]. interpolation [424, 51, 471, 472, 77, 173, 632, 290]. intersection [147]. Intertwining [580]. interval [440]. invariance [130]. Invariant [323, 509, 663, 582, 381, 574, 132, 475, 307, 576, 465, 337, 634, 350, 289, 422, 566]. Invariants [658, 693, 625]. inverse [58]. inversion [14]. invertibility [137]. invertible [389, 562]. involution [331, 121]. involving [276]. IP [30]. IP-sets [30]. Irrationalities [646, 698]. irreducible [414, 344]. irregularly [474]. Isometric [319, 321]. Isomorphism [627, 39, 693, 21]. Isomorphisms [271, 581, 37]. Isoperimetric [561]. isotropic [568]. iterated [370, 225]. iterative [77]. Iwahori [579]. Iwasawa [563].

Jackson [550]. Jacobi [256, 366, 434]. Jacobian [254, 631]. Joubert [640]. jump [361].

Kac [619]. Kähler [283]. kaleidoscope [397]. Kernel [294, 75, 573, 287, 522, 202]. kernels [130, 265]. Killing [634]. kind [300]. Kloosterman [210, 284]. knot [232]. Knotted [681]. Kodaira [629]. Königsberg [357]. Kostka [598]. Kottwitz [636]. Kukles [521].

lacunary [113]. Lagrange [471, 472, 632]. Laguerre [16, 131]. Langevin [171]. Langlands [671]. Lanteri [150]. Laplacian [697, 516, 574, 202]. Laplacians [244]. large [606]. Lashof [269]. Lattice [647, 386, 154, 119, 19]. lattices [73, 24, 292, 54, 397, 289]. law [225]. Lawrence [4]. laws [526, 534, 31]. Leeuw [377]. Leffler [600]. left [289]. Legendre [390, 630]. lemma [575, 417, 110]. lemme [110]. leur [245]. level [166, 577]. Liapunov [394]. Lie [179, 457, 582, 618, 448, 221, 15, 670, 260, 172, 42, 622, 633, 510, 374, 546, 532, 665, 96, 314, 352, 468, 581]. Lie-algebraic [633]. lifting [453]. lifts [498]. light [145]. Like [697, 638, 516]. Limit [677, 512, 523, 113, 526, 534, 483, 261, 286]. limiting [416]. limits [38, 338]. line [173]. linéaires [8, 226, 110]. Linear [686, 233, 367, 595, 63, 383, 410, 214, 271, 502, 226, 110]. links [223, 376]. L’invariant [634]. Liouville [596]. Lipschitz [626, 281, 165]. Littlewood [454, 351, 327]. Lizorkin [423]. Local [684, 145, 398, 521, 165, 111, 617, 470, 241, 8, 609, 453, 213, 615, 8]. Localization [621, 195]. locally [315, 463, 549, 559, 362, 198, 104, 501, 34]. Location [293]. Loeb [178]. logarithm [225]. logarithmes [226]. logarithms [226]. longest [286]. Loops [350, 196]. Lorentz [484, 351]. Lower [140, 192, 629, 363, 226]. Lusternik [615].

MacPherson [636]. Maddox [157]. Majorations [560]. manifold [148].

Manifolds [674, 514, 626, 52, 153, 473, 283, 600, 144, 82, 252, 569, 548]. Many [680]. map [494, 155]. Mappings [241, 656, 255, 158, 388, 429, 145, 27]. Maps [463, 549, 691, 644, 76, 588, 198, 191]. Marcinkiewicz [240]. marginals [415]. Markov [53, 496, 91, 361, 408]. Markusevic [298]. Martin [109]. matrices [140, 331, 256, 290, 443, 175]. matricial [156]. matrix [448, 371, 496, 538, 323]. maxima [108]. Maximal [644, 359, 330, 142, 185, 465, 489, 351, 97]. maximizations [560]. Mazur [267]. Mean [632, 471, 472, 54, 337]. means [347]. measurability [178]. measurable [558, 176]. Measure [295, 111, 209]. Measure-valued [295, 111]. Measures [215, 426, 535, 371, 415, 268, 340, 202]. medium [78]. Mellin [137, 421]. memoriam [217]. Meromorphic [642]. metabelian [170, 305]. metacirculant [627]. Method [671, 382, 204]. metric [451]. Metrics [437, 626, 244]. mild [151, 238]. Mills [162]. Minimal [89, 97, 334, 187, 223]. minimality [165]. minimax [146]. Minkowski [655]. Minorations [226]. Mittag [600]. Mittag-Leffler [600]. mixed [625, 322]. mixing [151, 238]. Moak [16]. Mod [261]. Model [654]. models [447, 2]. modifications [268]. Modular [680, 261, 26, 21, 489, 434]. module [539, 620, 309, 262]. modules [29, 608, 65, 609, 346, 613, 325, 146, 491]. moduli [9, 229]. modulo [92, 19]. modulus [287]. moment [538, 418]. moments [415, 528]. monochromatic [142]. monoid [614]. monoids [326]. monomial [231]. Monotone [125, 89]. monotonic [101]. Monotonicity [281, 366]. Moody [619]. Moore [328, 458, 441]. Mordell [541]. Morphisms [650]. Morse [597]. Motion [666, 462, 123]. Moufang [196, 653]. moved [410]. Moving [239, 404]. moyennables [622]. multi [93, 198, 490]. multi-dimensional [93, 490]. multi-valued [198]. Multidimensional [77]. multigraded [231]. Multiparameter [308, 557, 54]. multiple [553]. Multiplication [509, 122]. Multiplicities [677]. multiplier [131, 137, 327, 10]. Multipliers [369, 603, 240, 377, 617, 356, 228, 79]. multisigned [231]. multivalued [588, 191]. multivariable [523]. multivariate [167]. Murnaghan [579].

Nakayama [579]. nature [429]. Navier [129]. near [118]. near-best [118]. nearest [520]. Necessary [471, 472, 299, 131]. negative [629, 136]. neighborhoods [152]. Nets [652]. networks [265]. Neumann [697, 568, 255, 256, 516, 117, 79, 559]. neutral [304, 198, 323, 478, 224, 125]. Nevanlinna [678]. NFDEs [394]. nilpotence [20]. Nilpotency [665]. Nilpotent [598, 688, 121, 355, 552, 15, 614]. nilpotents [409]. Noguchi [429]. noise [287]. nombres [226]. Non [113, 676, 547, 469, 367, 127, 402, 629, 385, 410, 568, 536, 487]. non-connected [385]. non-defining [469]. Non-Gaussian [113, 127]. Non-Hermitian [547]. non-isotropic [568]. non-linear [367, 410]. non-negative [629]. non-polynomial [487]. non-singular [536]. non-smooth [402]. Non-Surjective [676]. Nonabelian [486, 76, 72, 163]. Noncommutative [585, 162]. Noncompact [649, 260, 123]. nondirecting [491]. nonexpansive [190]. nonincreasing [126]. nonlinear [208, 190, 276]. Nonlinearity [557]. Nonlocal [221]. nonnegative [91]. Nonnilpotent [56]. nonoscillation [304]. Nonsymmetric [291]. Norm [297, 391, 9, 229, 383, 573, 209, 13, 322, 555, 351]. norm-attaining [383]. Normal [507, 318, 205, 102, 124]. Normals [638]. norms [329, 589, 247]. Notes [435]. nuclear [23]. number [75, 119, 94, 223]. Numbers [683, 60, 18, 186, 226, 393]. numerator [324]. Numerical [492].

objects [58]. obstacles [456]. obstruction [243]. Occurrence [664]. occurs [48]. Odd [699, 599, 261]. ODE [477]. old [310]. One [540, 68, 83, 555, 351, 313]. one-dimensional [313]. One-parameter [68]. one-sided [555, 351]. open [362]. operations [245, 261, 335, 176, 245]. Operator [312, 505, 181, 391, 141, 52, 256, 416, 455, 351, 148, 97, 580]. Operators [699, 596, 89, 53, 90, 280, 573, 319, 473, 599, 389, 45, 26, 555, 299, 289, 339, 275, 607, 423]. optimization [266]. orbit [367, 53, 598, 552]. orbital [306, 59, 110]. orbitales [110]. order [88, 193, 526, 534, 627, 473, 372, 477, 478, 61, 49, 136, 234, 248, 224, 341, 98, 276]. Ordered [652, 142, 333, 235, 353]. orders [102]. orientable [463, 549]. oriented [569]. origin [191]. Orlicz [169]. Ornstein [127, 9, 229, 291, 225]. orthodox [73]. Orthogonal [268, 652, 545, 513, 436, 301, 371, 496, 517, 135, 528, 2, 523]. Oscillation [304, 478, 276]. Oscillations [224].

Packing [682]. packings [572]. Pair [643]. pairings [44]. Pairs [664, 323]. Pairwise [100]. Paley [327]. paper [150]. Parabolic [663, 544]. parallelizability [569]. Parameter [540, 294, 68]. Parametrizing [70]. Partial [641, 245, 498, 313]. partially [569]. particle [116, 520]. partielles [245]. parts [185]. path [294, 257]. paths [386]. patterns [48]. Peano [357]. Pelczyński [467]. Period [668, 103]. period-doubling [103]. periodic [493, 597, 148]. Periods [105, 521]. permutability [419]. permutation [122]. permutations [300, 345]. Perron [93]. perturbations [87]. Peterson [335]. Pexider [17]. Pick [678]. Pierce [212]. Pieri [584, 567]. pinching [247]. PL [514]. Planar [697, 516]. Planarity [292, 213]. Plane [683, 57]. pluri [195]. Point [647, 551, 103, 360, 190, 191, 176]. Points [108, 278, 402, 474, 169, 387, 334, 119, 61, 49, 136, 147, 188, 98]. pointwise [459]. polar [123]. polarized [629]. Poles [671, 645, 578, 642]. polycyclic [192, 543]. polydiscs [45]. polygon [602]. polyhedra [537]. Polynômes [218]. Polynomial [675, 448, 658, 132, 179, 331, 50, 590, 214, 487]. Polynomials [551, 250, 700, 513, 181, 138, 436, 329, 370, 293, 598, 300, 219, 194, 371, 496, 91, 268, 463, 133, 35, 135, 528, 118, 630, 16, 390, 60, 18, 186, 533, 523, 122, 366, 218]. polytopes [4, 119, 593, 399]. porous [78]. position [153]. Positive [138, 23, 74, 182, 540, 199, 479, 443, 437, 371, 389, 61, 49, 339, 175, 188, 98]. Positive-definite [479, 443]. positivity [408]. pour [570, 635, 696, 8, 115, 117, 560, 110]. power [460, 625, 428]. powers [574, 92, 511, 443]. Pragacz [386]. prescribed [324, 578, 250, 642]. preservation [104]. preserve [241]. preserving [255, 416, 408]. presheaves [621]. primary [62]. Prime [425, 60, 18, 595, 205, 586, 385, 186, 121]. prime-producing [186]. Primes [601, 680, 595, 460, 627, 438, 101, 407]. primitive [227, 221]. Primitivity [170, 305]. principle [484]. principles [191]. priori [446]. pro [387]. pro- [387]. probabilities [192]. probability [53]. problem [597, 627, 538, 117, 388, 47, 105, 120, 21, 489, 568, 357]. Problème [120, 117]. Problems [557, 88, 139, 243, 33, 199, 440]. processes [111, 506, 127, 9, 229, 294, 116, 404, 361, 54, 408, 291, 225]. producing [60, 18, 186]. Product [513, 470, 242, 627, 587, 428, 452, 231, 173, 56, 441, 189]. Products [669, 339, 141, 29, 586, 450, 406, 485, 336, 6, 364, 564, 273]. profinite [72, 163]. progressions [601]. projectionless [485]. projections [409]. Projective [674, 269, 231, 693, 492, 67, 464, 320, 346, 211, 94]. Prokhorov [451]. Proof [420, 550, 575]. Propagation [263]. Properties [638, 505, 294, 153, 383, 304, 428, 145, 466, 501, 71, 313, 150, 536]. property [7, 596, 53, 80, 467, 201]. Proto [27]. Proto-differentiation [27]. proximal [583, 277]. pseudo [152, 53, 69]. pseudo-actions [69]. pseudo-orbit [53]. pseudoconvex [388, 607]. Purely [587, 340].

Quadrangles [653]. Quadratic [695, 646, 698, 139, 1, 60, 18, 186, 420, 531, 236, 124]. quadrics [536]. Quantization [552]. Quantizations [262]. Quantized [512]. quantum [343, 390, 563, 288, 504]. Quartic [207, 683, 1, 124]. Quasi [356, 375, 480, 515, 629, 639, 639]. Quasi-commutative [639]. quasi-commutatives [639]. quasi-heredity [480]. quasi-homogeneous [515]. Quasi-multipliers [356]. quasi-polarized [629]. Quasi-regular [375]. quasibundles [432]. Quasiconformal [675]. quasiconvex [499]. questions [538]. queues [396]. Quillen [305]. Quiver [650]. Quotient [661, 641, 414]. Quotients [649, 457, 559].

R. [96]. radially [78]. Radical [663, 571]. radius [419]. Ramanujan [412, 92, 32]. ramified [539, 486]. Ramsey [201]. Random [228, 138]. Randomization [229]. range [558, 392, 137, 87]. ranges [156]. rank [305, 387, 541, 264, 518, 338, 10]. ranks [296]. rates [451]. Rational [424, 427, 211, 545, 548, 268, 324, 578, 556, 567]. Ray [556]. Razumikhin [394]. Real [320, 139, 166, 358, 359, 330, 628, 549, 323, 60, 18, 186, 173, 363, 420, 160, 400, 188, 122, 10]. realization [613]. Realizations [691]. rearrangement [475]. rearrangement-invariant [475]. rebroussements [396]. Reconstruction [474]. recovery [449]. rectifiable [411]. reduced [186, 336, 521]. Reducibility [385]. Reduction [638, 338, 417, 19]. reductive [193, 614]. Rees [149]. refinable [553]. Reflection [686, 687, 130, 579, 580]. Regular [691, 178, 514, 610, 593, 95, 249, 375]. Related [682, 74, 293, 146]. relating [395]. relation [245]. relationship [245]. Relative [403, 210, 284, 366]. Remarks [109]. renormalization [367, 611]. reordered [56]. Repercussions [47]. replicable [427]. represent [282]. Representation [64, 673, 197, 631, 511, 504]. Representations [676, 650, 343, 659, 260, 672, 96, 23, 414, 344, 231, 320, 542, 590, 42, 398, 160, 2, 623, 87]. residual [500]. Residue [421, 186, 616]. Résolubilité [8]. resolutions [180]. Resonance [699, 599, 11]. resonances [456]. respect [371]. Restricted [374, 34, 197]. Restricting [42]. restriction [498]. restrictions [292]. result [439]. results [597, 242, 395]. reversible [520]. Riccati [547]. ridge [290]. Riemann [594, 256, 600, 55, 161, 605]. Riemannian [181, 600, 244]. Riesz [179, 450, 622, 482]. rigid [628]. rigidity [388]. Ring [673, 571, 166, 158, 395, 50, 591, 543, 488, 504]. Rings [657, 465, 380, 282, 219, 495, 586, 527, 372, 543, 121, 431, 94, 271, 86, 55, 21]. Riordan [519]. Robinson [217]. Rogers [412]. Root [692, 582, 551, 397]. roots [542]. rotation [360]. rough [573]. rules [579, 476, 173]. run [48, 286].

Samelson [6]. sample [474]. Satisfying [680]. scalaires [115]. scalar [437, 115]. Scattering [87]. scheme [309]. Schemes [652, 446]. Schensted [2]. Schilder [462]. Schnirelmann [615]. Schoenberg [416].

Schreier [532]. Schrödinger [699, 148, 599]. Schur [177, 517, 94]. Second [248, 300, 477, 478, 276]. Second-order [248]. sections [619, 486, 350, 160]. Segal [670, 171]. segments [425, 601, 503]. selections [168]. self [6]. self-homotopy [6]. Semi [689, 8, 440]. Semi-Affine [689]. semi-infinite [440]. semi-linéaires [8]. semidirect [242]. semiflows [125]. Semigroup [657, 289]. Semigroups [540, 73, 585, 405, 58, 479, 190, 353, 176]. Semilattices [669]. semilinear [8, 102, 263, 199, 108]. semimartingales [46]. semisimple [618]. sens [117]. sense [117]. Sensitivity [277]. Separable [638, 84]. Separating [334]. separation [633]. Sequence [151, 114, 157, 651, 238]. sequences [505, 182, 429, 556]. Sergeev [386]. Series [645, 113, 331, 132, 28, 529, 592, 564, 236, 85, 327]. set [383, 183, 101, 337, 44, 27]. set-valued [183, 27]. Sets [558, 682, 30, 89, 142, 81, 577, 349, 411]. several [391, 392, 216]. Severi [13]. shadowing [53]. Shahidi [671]. Shape [588, 572, 416]. Sharpness [364]. shift [319, 422]. shift-invariant [422]. short [601, 87]. sided [555, 351]. Siegel [645, 568]. Sierpinski [666, 682]. Signed [641]. signs [315]. Similar [694]. Similarity [692]. Simple [658, 485, 359, 330, 587, 260, 407, 548]. simplicial [621]. Simplicity [336]. Simply [38, 487]. single [477]. singular [295, 602, 287, 440, 536]. Singularities [317, 477, 263]. Site [663]. Six [595, 196]. size [200]. sizes [100]. Skeleton [159]. skew [425, 386, 50, 323, 121, 265]. Slicings [168]. Smale [442]. small [89, 489]. Smooth [662, 616, 402]. smoothness [228, 603]. Sobolev [426, 461, 508, 454]. solenoids [470]. solid [12]. solitons [624]. solution [568]. solutions [88, 624, 493, 597, 459, 258, 117, 78, 477, 199, 547, 313, 544]. Solvability [440, 8, 108]. solvable [497, 42, 95, 249]. Some [412, 589, 305, 529, 682, 236, 665, 700, 203, 536, 446, 280, 294, 208, 1, 79, 388, 43, 577, 519, 48, 440, 533, 189, 150]. sous [583, 546]. sous-algèbres [546]. sous-différentiel [583]. Space [690, 644, 368, 437, 53, 381, 66, 464, 185, 171, 81, 46, 211, 291, 548, 123, 191]. Space-Forms [690]. spaced [474]. Spaces [677, 638, 660, 540, 661, 509, 512, 342, 178, 445, 180, 328, 594, 181, 369, 141, 617, 525, 114, 426, 535, 625, 461, 303, 220, 484, 157, 169, 209, 604, 270, 475, 333, 508, 429, 476, 198, 630, 602, 261, 321, 26, 233, 622, 454, 466, 248, 6, 482, 235, 442, 467, 565, 175, 503, 423]. Spacetime [655]. sparse [185]. Spatial [506, 376]. special [341]. specialised [274]. Species [654]. Spectra [452, 458, 206, 526, 534]. Spectral [240, 649, 516, 697, 608, 419, 188, 87]. spectrum [139, 166, 301, 256, 68, 500, 148]. Spheres [513]. spherical [437, 108]. spline [63]. Square [671, 234]. square-free [234]. squared [9, 229]. squares [639]. stability [348]. stabilized [139]. Stable [668, 569, 458, 143, 422, 99]. stably [485]. standard [417]. star [401, 593]. star-polytopes [593]. states [188]. Steenrod [639]. Stein [7]. Steinberg [457]. Steiner [301]. Steps [684]. Stiefel [674]. Stirling [393]. Stochastic [46, 171, 313]. Stokes [129]. stratification [149]. strictly [456]. strips [40]. Strong [114, 660]. Strongly [80, 152]. Structural [466]. Structure [95, 249, 470, 177, 539, 620, 157, 349, 10, 99]. structures [328, 187, 237]. Struppa [150]. stunted [211]. sub [244]. sub-Laplacians [244]. Subalgebras [563, 448, 360, 159, 546, 532]. subcategories [487]. subdifferential [583, 476]. Subdifferentials [648]. Subgradient [281, 27]. subgroup [378, 75]. Subgroups [571, 663, 62, 279, 221, 495, 102, 499, 348, 70]. subharmonic [195]. sublattice [119]. Sublattices [694]. Submodules [692]. submultiplicativity [419]. subnormal [455]. subordination [506]. Subquotients [687]. Subregular [688]. subrings [465]. subsemigroups [362]. Subspaces [475, 509, 381, 559, 576, 323, 651, 565, 422]. subvarities [536]. sufficient [471, 472, 299]. Sufficiently [502]. sum [515, 81]. summability [204]. summands [143, 491]. Summation [668, 1, 529, 236]. Summing [531]. sums [40, 71, 325]. super [29, 462]. super-Brownian [462]. super-Virasoro [29]. superadditive [404, 54]. supercritical [116]. supercuspidal [398]. Superresolution [451]. supersingular [438]. supersymmetric [386]. supported [63, 66]. Supporting [638]. surfaces [594, 256, 629, 600, 259, 463, 549, 464, 561, 396, 55, 161, 605]. Surjective [676, 258]. surjectivité [245]. surjectivity [245]. Suslin [305]. Sweeping [505]. Sylow [75]. symbols [258, 275]. Symmetric [690, 651, 106, 181, 525, 220, 549, 78, 333, 323, 11, 233, 123]. Symmetries [206, 83]. symplectic [317, 517, 162]. Synergy [257]. system [116, 521]. Systems [692, 139, 582, 301, 258, 604, 418, 577, 520, 531, 273, 125, 277].

Tableaux [613]. tactics [267]. TAF [311, 272]. tame [75, 487]. Tauberian [347, 310, 204]. Tauberian- [347]. Taylor [132]. technique [69, 532]. temperatures [604]. Tempered [66, 623]. Tensor [141, 29, 669, 672, 564, 452, 406, 262, 2, 253]. tent [423]. Tessellations [690]. Testing [663]. th [88]. Thaine [382]. their [23, 470, 558, 294, 66, 168, 245, 498, 350, 421, 374, 10]. Theorem [655, 640, 112, 183, 496, 462, 497, 79, 333, 93, 198, 499, 54, 46, 520, 189, 190, 204]. theorems [377, 379, 414, 404, 600, 305, 394, 529, 308, 429, 347, 348, 454, 310, 176, 327]. theoretic [483, 67, 536]. theories [257, 621]. Theory [697, 679, 688, 663, 179, 391, 154, 243, 516, 628, 132, 134, 395, 173, 160, 107, 297, 615, 162, 353, 97, 87]. theta [254, 623]. third [307]. three [100, 129, 95, 249]. tic [400]. till [48]. time [138, 294, 48]. times [111]. Toeplitz [275]. top [342]. Topological [383, 429, 198, 150]. Topologizing [358]. topology [432]. tordues [110]. Tori [686, 359, 330, 162]. Toroidal [659, 691]. torsion [481, 317, 584]. torus [576]. Total [12]. Totally [628]. Towers [649]. Trace [619, 543, 605]. Traces [676]. Tractable [637]. Transfer [668, 193]. Transferability [669]. Transform [670, 22, 384, 322, 14]. transformation [367, 1, 37]. transformations [611, 260, 529, 233, 410, 592, 251]. transformées [622]. transforms [179, 66, 622, 603, 137, 421, 297]. transitions [523]. translates [63]. transversality [198]. Triebel [423]. trigonometric [113]. Trimèche [617]. triple [301]. truncated [538]. tubular [152]. Twisted [380, 541, 231, 59, 441, 110, 504]. Two [209, 35, 313, 342, 515, 294, 429, 81, 555, 120, 577, 536]. two-time [294]. Two-weight [209]. Type [655, 670, 104, 653, 328, 377, 493, 426, 91, 462, 64, 416, 452, 394, 78, 171, 68, 50, 555, 454, 55, 523, 607, 123, 581]. types [211, 31, 353]. typical [3].

Uhlenbeck [127, 9, 229, 291, 225]. Ulam [47]. ultraproducts [373]. ultraspherical [630]. unconditional [565]. Unification [355]. uniform [112, 258, 41]. Unipotent [306, 663]. Uniqueness [654]. Unit [545, 417, 591, 86]. unitals [57]. Unitary [687, 457, 29, 575]. Units [586, 372, 380, 382, 196, 289]. unity [542]. Univalent [656, 158]. Universal [24, 178, 589, 353]. Upper [599, 699, 612]. uscos [89]. valeurs [218]. valuation [65]. valuations [395]. Value [128, 440]. valued [138, 111, 127, 183, 303, 295, 198, 477, 27, 434]. Values [646, 698, 412, 218, 242, 129, 40]. vanishing [158]. variable [391]. variables [138, 392, 120, 633, 216]. variation [206]. Variational [557]. varieties [354, 355, 368, 598, 618, 24, 13, 104, 80, 58, 411]. variety [631, 614]. vector [515, 303, 81, 434, 253]. vector-valued [303]. Verma [613]. Veronesean [259]. versality [152]. Vertical [62]. very [492]. via [322, 291, 567, 277]. view [243]. view-obstruction [243]. Vilenkin [327]. Virasoro [29, 659]. Virasoro-Toroidal [659]. Vitali [112]. Voronoi [397].

waiting [48]. Waldspurger [439]. Ward [624]. wave [287, 87]. Weak [401, 377, 466, 289]. Weak-star [401]. Weakly [346, 340, 388]. wedge [143]. Weight [513, 209, 135, 555, 410, 397]. Weighted [126, 51, 573, 660, 222, 322, 351, 239, 461, 209, 308, 482, 202]. Weights [106, 471, 472, 632]. Weil [541]. Weiss [7]. Weyl [79]. which [278, 241, 558, 563, 56, 2, 201]. Whitehead [302]. whose [380]. width [213]. Wielandt [378]. wildly [539]. Wilson [35]. wise [246]. without [190]. Witt [282, 527, 497, 395, 262, 431, 634]. Wold [672]. word [499].

Yang [162]. Young [40, 322, 364].

Zero [545, 68, 250, 10]. Zeros [370, 700, 293, 118, 642, 533]. zonal [463].

References

Gasper:1990:IBS

[1] George Gasper and Mizan Rahman. An indefinite bibasic summation formula and some quadratic, cubic and quartic summation and transformation formulas. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):1–27, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Proctor:1990:SAW

[2] Robert A. Proctor. A Schensted algorithm which models tensor representations of the orthogonal group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 28–49, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barany:1990:DTC

[3] Imre Bárány and Tudor Zamfirescu. Diameters in typical convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 50–61, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bayer:1990:LP

[4] Margaret Bayer and Bernd Sturmfels. Lawrence polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):62–79, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sankaran:1990:ACH

[5] P. Sankaran and K. Varadarajan. Acyclicity of certain homeomorphism groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):80–94, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moller:1990:SPS

[6] Jesper Michael Møller. Samelson products in spaces of self-homotopy equivalences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):95–108, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asmar:1990:SWP

[7] Nakhlé Asmar. On a Stein and Weiss property of the conjugate function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):109–125, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dehman:1990:RLP

[8] B. Dehman. Résolubilité local pour des équations semi-linéaires complexes. (French) [Local solvability of semilinear complex equations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):126–140, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Csorgo:1990:MCG

[9] Miklós Csörgö and Zheng Yan Lin. On moduli of continuity for Gaussian and l^{2}-norm squared processes generated by Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 141–158, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:1990:ARR

[10] Shuang Zhang. C^{∗}-algebras with real rank zero and the internal structure of their corona and multiplier algebras. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):159–190, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:1990:CSH

[11] Cheng Zhi Li and Christiane Rousseau. Codimension 2 symmetric homoclinic bifurcations and application to 1: 2 resonance. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):191–212, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borger:1990:TCS

[12] Reinhard Börger and Walter Tholen. Total categories and solid functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):213–229, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kang:1990:CBS

[13] Ming Chang Kang. Constructions of Brauer–Severi varieties and norm hypersurfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):230–238, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1990:DHT

[14] O. P. Singh and J. N. Pandey. The ndimensional Hilbert transform of distributions, its inversion and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):239–258, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drensky:1990:AFN

[15] Vesselin Drensky and C. K. Gupta. Automorphisms of free nilpotent Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):259–279, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koekoek:1990:GML

[16] Roelof Koekoek. A generalization of Moak’s q-laguerre polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 280–303, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deeba:1990:PFE

[17] E. Y. Deeba and E. L. Koh. The Pexider functional equations in distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):304–314, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Louboutin:1990:PPQ

[18] Stéphane Louboutin. Prime producing quadratic polynomials and classnumbers of real quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 315–341, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Symonds:1990:RLM

[19] Peter Symonds. The reduction of an RG-lattice modulo p^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):342–364, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kearnes:1990:AN

[20] Keith A. Kearnes. Atomicity and nilpotence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):365–382, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rohl:1990:ACA

[21] Frank Röhl. On automorphisms of complete algebras and the isomorphism problem for modular group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 383–394, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feichtinger:1990:GAA

[22] Hans G. Feichtinger. Generalized amalgams, with applications to Fourier transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):395–409, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Belanger:1990:PFN

[23] Alain Belanger and Erik G. F. Thomas. Positive forms on nuclear ∗-algebras and their integral representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 410–469, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goralcik:1990:UVL

[24] P. Goralcík, V. Koubek, and J. Sichler. Universal varieties of (0, 1)-lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):470–490, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parry:1990:BBF

[25] Charles J. Parry. Bicyclic bicubic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):491–507, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labuda:1990:DAO

[26] Iwo Labuda. Disjointly additive operators and modular spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 508–519, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poliquin:1990:PDS

[27] René A. Poliquin. Proto-differentiation of subgradient set-valued mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):520–532, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hatcher:1990:HS

[28] Rhonda Lee Hatcher. Heights and L-series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):533–560, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bremner:1990:TPU

[29] Murray R. Bremner. Tensor products of unitary super-Virasoro modules with central charge 7/10. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):561–574, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berend:1990:ISC

[30] Daniel Berend. IP-sets on the circle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):575–589, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Truss:1990:FCL

[31] J. K. Truss. The failure of cancellation laws for equidecomposability types. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):590–606, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zeng:1990:FRB

[32] Jiang Zeng. Sur une formule de Ramanujan–Bailey. (French) [On a formula of Ramanujan–Bailey]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 607–618, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hengartner:1990:EPC

[33] Walter Hengartner and Wojciech Szapiel. Extremal problems for the classes S_{R}^{−p} and T_{R}^{−p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):619–645, ????

1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snaith:1990:RDH

[34] Victor Snaith. Restricted determinantal homomorphisms and locally free class groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):646–658, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1990:TFA

[35] M. E. H. Ismail, J. Letessier, G. Valent, and J. Wimp. Two families of associated Wilson polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):659–695, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:1990:FED

[36] John A. Baker. Functional equations, distributions and approximate identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):696–708, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamoureux:1990:AIT

[37] Michael P. Lamoureux. Analytic isomorphisms of transformation group C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):709–730, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pare:1990:SCL

[38] Robert Paré. Simply connected limits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):731–746, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kwak:1990:ICG

[39] Jin Ho Kwak and Jaeun Lee. Isomorphism classes of graph bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 747–762, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:1990:EVD

[40] I. P. Goulden. Exact values for degree sums over strips of Young diagrams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):763–775, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakazi:1990:CDU

[41] Takahiko Nakazi. Commuting dilations and uniform algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):776–789, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lipsman:1990:RRC

[42] R. L. Lipsman. Restricting representations of completely solvable Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):790–824, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kiming:1990:ECS

[43] Ian Kiming. Explicit classifications of some 2-extensions of a field of characteristic different from 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):825–855, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oda:1990:HSA

[44] Nobuyuki Oda. The homotopy set of the axes of pairings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):856–868, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jafari:1990:BCC

[45] F. Jafari. On bounded and compact composition operators in polydiscs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):869–889, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leon:1990:SFT

[46] Jorge A. León. Stochastic Fubini theorem for semimartingales in Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):890–901, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Just:1990:RPE

[47] Winfried Just. Repercussions of a problem of Erdős and Ulam on density ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):902–914, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mori:1990:MWT

[48] Tamás F. Móri. More on the waiting time till each of some given patterns occurs as a run. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):915–932, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1990:PIB

[49] Lee Lorch and Peter Szego. On the points of inflection of Bessel functions of positive order. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(5):933–948, October 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigenda [61].

Kikumasa:1990:ACS

[50] Isao Kikumasa. Automorphisms of a certain skew polynomial ring of derivation type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):949–958, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1990:WII

[51] R. C. Brown and D. B. Hinton. Weighted interpolation inequalities and embeddings in R^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):959–980, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DAtri:1990:ECO

[52] J. E. D’Atri and I. Dotti Miatello. Eigenvalues of the curvature operator for certain homogeneous manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 981–999, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyarsky:1990:POS

[53] Abraham Boyarsky and Pawel Góra. The pseudo-orbit shadowing property for Markov operators in the space of probability density functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 1000–1017, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1990:MET

[54] Felix Lee. A mean ergodic theorem for multiparameter superadditive processes on Banach lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):1018–1040, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ripoll:1990:RCT

[55] Pascual Cutillas Ripoll. On rings with a certain type of factorization and compact Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 1041–1052, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maj:1990:NGW

[56] M. Maj and S. E. Stonehewer. Nonnilpotent groups in which every product of four elements can be reordered. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 1053–1066, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kestenband:1990:FUH

[57] Barbu C. Kestenband. A family of unitals in the Hughes plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??): 1067–1083, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Margolis:1990:FOC

[58] S. W. Margolis, J. C. Meakin, and J. B. Stephen. Free objects in certain varieties of inverse semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):1084–1097, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Joyner:1990:TOI

[59] David Joyner. On twisted orbital integral identities for PGL(3) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):1098–1130, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Louboutin:1990:APP

[60] S. Louboutin. Addendum: “Prime producing quadratic polynomials and class-numbers of real quadratic fields”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42 (??):1131, ???? 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1990:CPI

[61] Lee Lorch and Peter Szego. Corrigenda to: “On the points of inflection of Bessel functions of positive order. I”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(6): 1132, December 1990. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [49].

Benabdallah:1991:VSP

[62] K. Benabdallah, B. Charles, and A. Mader. Vertical subgroups of primary abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):3–18, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chui:1991:CBS

[63] Charles K. Chui and Amos Ron. On the convolution of a box spline with a compactly supported distribution: linear independence for the integer translates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):19–33, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frappier:1991:RFI

[64] Clément Frappier. Representation formulas for integrable and entire functions of exponential type. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):34–47, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fuchs:1991:BMV

[65] L. Fuchs and E. Monari Martinez. Butler modules over valuation domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 48–60, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gabardo:1991:TDS

[66] Jean-Pierre Gabardo. Tempered distributions supported on a half-space of R^{n} and their Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 61–88, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gottlieb:1991:GTA

[67] Alex D. Gottlieb and Joseph Lipman. Group-theoretic axioms for projective geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):89–107, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kawahigashi:1991:OPA

[68] Yasuyuki Kawahigashi. One-parameter automorphism groups of the injective factor of type II_{1} with Connes spectrum zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):108–118, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Loring:1991:BTP

[69] Terry A. Loring. Berg’s technique for pseudo-actions with applications to AF embeddings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):119–157, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maclachlan:1991:PFS

[70] C. Maclachlan and A. W. Reid. Parametrizing Fuchsian subgroups of the Bianchi groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):158–181, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oskolkov:1991:FPI

[71] K. I. Oskolkov. On functional properties of incomplete Gaussian sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 182–212, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ulbrich:1991:NPG

[72] K.-H. Ulbrich. On nonabelian H^{2} for profinite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):213–224, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Auinger:1991:BCL

[73] Karl Auinger. Boolean congruence lattices of orthodox semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 225–241, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bloom:1991:PDR

[74] Walter R. Bloom and Paul Ressel. Positive definite and related functions on hypergroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):242–254, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauckmann:1991:SST

[75] Boris Brauckmann. The 2-sylow-subgroup of the tame kernel of number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):255–264, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullejos:1991:DNC

[76] Manuel Bullejos and Antonio M. Cegarra. A 3-dimensional nonabelian cohomology of groups with applications to homotopy classification of continuous maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):265–296, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deslauriers:1991:MII

[77] Gilles Deslauriers, Jacques Dubois, and Serge Dubuc. Multidimensional iterative interpolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):297–312, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hernandez:1991:HER

[78] Gastón E. Hernández and Ioannis M. Roussos. The Hölder exponent for radially symmetric solutions of porous medium type equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):313–321, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higson:1991:WNT

[79] Nigel Higson and Mikael Rørdam. The Weyl–von Neumann theorem for multipliers of some AF-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 322–330, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kiss:1991:SAV

[80] E. Kiss and M. Valeriote. Strongly abelian varieties and the Hamiltonian property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):331–346, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krantz:1991:VST

[81] Steven G. Krantz and Harold R. Parks. On the vector sum of two convex sets in space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):347–355, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreck:1991:CMF

[82] Matthias Kreck and Georgia Triantafillou. On the classification of manifolds up to finite ambiguity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):356–370, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalliongis:1991:SGO

[83] John Kalliongis and Andy Miller. The symmetries of genus one handlebodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):371–404, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Navarro:1991:FCC

[84] Gabriel Navarro. Fong characters and correspondences in pi-separable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):405–412, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramsamujh:1991:CED

[85] T. I. Ramsamujh. The complexity of everywhere divergent Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 413–424, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Riley:1991:GRH

[86] David M. Riley. Group rings with hypercentral unit groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):425–434, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yamamoto:1991:STS

[87] Kazuhiro Yamamoto. Scattering theory and spectral representations for general wave equations with short range perturbations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):435–448, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1991:DSO

[88] W. Allegretto and L. S. Yu. Decaying solutions of 2mth order elliptic problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):449–460, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1991:MCU

[89] Jonathan Borwein, Simon Fitzpatrick, and Petàr Kenderov. Minimal convex uscos and monotone operators on small sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(3):461–476, June 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bradley:1991:IO

[90] R. E. Bradley. On induced operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):477–494, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdelyi:1991:BMT

[91] Tamás Erdélyi. Bernstein and Markov type inequalities for generalized nonnegative polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):495–505, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1991:RCM

[92] Kim Hughes. Ramanujan congruences for p_{−k}(n) modulo powers of 17. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 506–525, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jurkat:1991:CMD

[93] W. B. Jurkat and R. W. Knizia. A characterization of multi-dimensional Perron integrals and the fundamental theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):526–539, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nelis:1991:SPS

[94] Peter Nelis. Schur and projective Schur groups of number rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):540–558, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ninomiya:1991:SSG

[95] Yasushi Ninomiya. Structure of psolvable groups with three p-regular classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):559–579, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [249].

Strade:1991:RDL

[96] Helmut Strade. Representations of the (p^{2} − 1)-dimensional Lie algebras of R. E. Block. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):580–616, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1991:MMO

[97] M. W. Wong. Minimal and maximal operator theory with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 617–627, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1991:PIB

[98] R. Wong and T. Lang. On the points of inflection of Bessel functions of positive order. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):628–651, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:1991:SSC

[99] Ying Bo Zhang. The structure of stable components. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):652–672, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Colbourn:1991:PBD

[100] Charles J. Colbourn, Alexander Rosa, and Douglas R. Stinson. Pairwise balanced designs with block sizes three and four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):673–704, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koninck:1991:AFM

[101] Jean-Marie De Koninck, Imre Kátai, and Armel Mercier. Additive functions monotonic on the set of primes. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):705–720, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Droste:1991:AGH

[102] M. Droste, W. C. Holland, and H. D. Macpherson. Automorphism groups of homogeneous semilinear orders: normal subgroups and commutators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 721–737, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erbe:1991:EFP

[103] L. H. Erbe, K. Geba, and W. Krawcewicz. Equivariant fixed point index and the period-doubling cascades. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):738–747, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kearnes:1991:TPL

[104] Keith A. Kearnes. Type preservation in locally finite varieties with the CEP. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):748–769, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lichtin:1991:PAD

[105] Ben Lichtin. Periods and the asymptotics of a Diophantine problem. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):770–791, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Michler:1991:WCG

[106] G. O. Michler and J. B. Olsson. Weights for covering groups of symmetric and alternating groups, p ≠ 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(4): 792–813, August 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Piacenza:1991:HTD

[107] Robert J. Piacenza. Homotopy theory of diagrams and CW-complexes over a category. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):814–824, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schechter:1991:PSM

[108] Martin Schechter and Kyril Tintarev. Points of spherical maxima and solvability of semilinear elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):825–831, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Todorcevic:1991:RMA

[109] Stevo Todorcević. Remarks on Martin’s axiom and the continuum hypothesis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):832–851, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Waldspurger:1991:IOT

[110] J.-L. Waldspurger. Sur les intégrales orbitales tordues pour les groupes linéaires: un lemme fondamental. (French) [On twisted orbital integrals for linear groups: a fundamental lemma]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43 (??):852–896, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barlow:1991:CLT

[111] Martin T. Barlow, Steven N. Evans, and Edwin A. Perkins. Collision local times and measure-valued processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 897–938, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bator:1991:VIC

[112] Elizabeth M. Bator, Russell G. Bilyeu, and Paul W. Lewis. The Vitali integral convergence theorem and uniform absolute continuity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):939–947, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berkes:1991:NGL

[113] I. Berkes. Non-Gaussian limit distributions of lacunary trigonometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):948–959, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buntinas:1991:SBS

[114] Martin Buntinas and Naza Tanović-Miller. Strong boundedness and strong convergence in sequence spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 960–974, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferland:1991:FPE

[115] René Ferland. Fluctuations pour des équations de Boltzmann scalaires. (French) [Fluctuations for scalar Boltzmann equations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):975–984, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fernandez:1991:CCG

[116] Begoña Fernández and Luis G. Gorostiza. A criterion of convergence of generalized processes and an application to a supercritical branching particle system. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):985–997, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guennoun:1991:ESA

[117] Zine E. A. Guennoun. Existence de solutions au sens de Carathéodory pour le problème de Neumann y′′ = f(t, y, y′). (French) [Existence of solutions in the Carathéodory sense for the Neumann problem y′′ = f(t, y, y′)]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 998–1009, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ivanov:1991:BZP

[118] K. G. Ivanov, E. B. Saff, and V. Totik. On the behavior of zeros of polynomials of best and near-best approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1010–1021, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lagarias:1991:BLP

[119] Jeffrey C. Lagarias and Günter M. Ziegler. Bounds for lattice polytopes containing a fixed number of interior points in a sublattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1022–1035, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Madi:1991:PCD

[120] Nour Saïd Madi. Problème de Cauchy à deux variables fuchsiennes.

(French) [Cauchy problem in two Fuchsian variables]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1036–1044, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martindale:1991:NID

[121] W. S. Martindale III and C. Robert Miers. Nilpotent inner derivations of the skew elements of prime rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1045–1054, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tautz:1991:EHC

[122] Walter Tautz, Jaap Top, and Alain Verberkmoes. Explicit hyperelliptic curves with real multiplication and permutation polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1055–1064, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1991:BMS

[123] J. C. Taylor. Brownian motion on a symmetric space of noncompact type: asymptotic behaviour in polar coordinates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1065–1085, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vaughan:1991:NCQ

[124] Theresa P. Vaughan. The normal closure of a quadratic extension of a cyclic quartic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1086–1097, ????

1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wu:1991:MSG

[125] Jian Hong Wu and H. I. Freedman. Monotone semiflows generated by neutral functional-differential equations with application to compartmental systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1098–1120, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andersen:1991:WGH

[126] Kenneth F. Andersen. Weighted generalized Hardy inequalities for nonincreasing functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1121–1135, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bojdecki:1991:GNG

[127] Tomasz Bojdecki and Luis G. Gorostiza. Gaussian and non-Gaussian distribution-valued Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1136–1149, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1991:VGD

[128] H.-H. Brungs and J. Gräter. Value groups and distributivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1150– 1160, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1991:NSD

[129] G. F. D. Duff. Navier–Stokes derivative estimates in three dimensions with boundary values and body forces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 1161–1212, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dunkl:1991:IKR

[130] Charles F. Dunkl. Integral kernels with reflection group invariance. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 1213–1227, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gasper:1991:NMC

[131] George Gasper and Walter Trebels. On necessary multiplier conditions for Laguerre expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1228–1242, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1991:PIT

[132] John E. Gilbert. Polynomial invariant theory and Taylor series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 1243–1262, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1991:ACH

[133] Dharma P. Gupta, Mourad E. H. Ismail, and David R. Masson. Associated continuous Hahn polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):

1263–1280, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hansen:1991:GCG

[134] Vagn Lundsgaard Hansen and Peter V. Petersen. Groups, coverings and Galois theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1281–1293, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1991:CWF

[135] Mourad E. H. Ismail, David R. Masson, and Mizan Rahman. Complex weight functions for classical orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1294–1308, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1991:IPB

[136] Lee Lorch, Martin E. Muldoon, and Peter Szego. Inflection points of Bessel functions of negative order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??): 1309–1322, ???? 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McBride:1991:RIC

[137] A. C. McBride and W. J. Spratt. On the range and invertibility of a class of Mellin multiplier transforms. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(6):1323– 1338, December 1991. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). To Professor Tim Rooney with best wishes.

Baker:1992:PPT

[138] B. M. Baker and D. E. Handelman. Positive polynomials and time dependent integer-valued random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1): 3–41, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barkwell:1992:GSS

[139] Lawrence Barkwell, Peter Lancaster, and Alexander S. Markus. Gyroscopically stabilized systems: a class of quadratic eigenvalue problems with real spectrum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):42–53, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:1992:LBM

[140] Grahame Bennett. Lower bounds for matrices. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):54–74, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blecher:1992:TPO

[141] David P. Blecher. Tensor products of operator spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):75–90, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duffus:1992:COS

[142] D. Duffus, V. Rödl, N. Sauer, and R. Woodrow. Coloring ordered sets to avoid monochromatic maximal chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (1):91–103, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:1992:CSW

[143] John C. Harris. On certain stable wedge summands of B(Z/p)_{+}^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1): 104–118, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hempel:1992:HBC

[144] John Hempel. Homology of branched coverings of 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):119– 134, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lyzzaik:1992:LPL

[145] Abdallah Lyzzaik. Local properties of light harmonic mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1): 135–153, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudlof:1992:MRM

[146] Peter Rudlof. On minimax and related modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):154–166, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sodhi:1992:CPH

[147] Amar Sodhi. The conductor of points having the Hilbert function of a complete intersection in P^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):167– 179, February 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sunada:1992:GAS

[148] Toshikazu Sunada. Group C^{∗}-algebras and the spectrum of a periodic schrödinger operator on a manifold. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (1):180–193, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:1992:SGA

[149] Ti Wang. A stratification given by Artin–Rees estimates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):194–205, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wisniewski:1992:TPS

[150] Jarosław A. Wiśniewski. On topological properties of some coverings. an addendum to a paper of A. Lanteri and D. C. Struppa: “Topological properties of cyclic coverings branched along an ample divisor”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(1):206–214, February 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:1992:SEM

[151] Qing Zhang. Sequence entropy and mild mixing. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):215–224, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akahori:1992:CVD

[152] Takao Akahori. A criterion for versality of deformations of tubular neighborhoods of strongly pseudo convex boundaries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):225–233, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daverman:1992:GPP

[153] R. J. Daverman and D. Repovs. General position properties that characterize 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):234–251, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1992:DCL

[154] Alan Day. Doubling constructions in lattice theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):252–269, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duszak:1992:DAM

[155] Zbigniew Duszak. On the degree of an analytic map germ. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):270–279, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farenick:1992:CMR

[156] D. R. Farenick. C^{∗}-convexity and matricial ranges. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):280–297, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grosse-Erdmann:1992:SSS

[157] Karl-Goswin Grosse-Erdmann. The structure of the sequence spaces of Maddox. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):298–307, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hengartner:1992:UHR

[158] Walter Hengartner and Jan Szynal. Univalent harmonic ring mappings vanishing on the interior boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 308–323, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:1992:SS

[159] Hua Xin Lin. Skeleton C^{∗}-subalgebras.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):324–341, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pfaffenberger:1992:CGT

[160] W. E. Pfaffenberger and J. Phillips.

Commutative Gel’fand theory for real Banach algebras: representations as sections of bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):342–356, ????

1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sakai:1992:CRS

[161] Makoto Sakai. Continuations of Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):357–367, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spera:1992:SAY

[162] Mauro Spera. A symplectic approach to Yang–Mills theory for noncommutative tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):368–387, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ulbrich:1992:NPG

[163] K.-H. Ulbrich. On nonabelian H^{2} for profinite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):388–399, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:1992:BCC

[164] Xiaolu Wang. A bivariant Chern character. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):400–435, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zeidan:1992:LML

[165] Vera Zeidan. Local minimality of a Lipschitz extremal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):436–448, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barton:1992:RSH

[166] Susan Maureen Barton. The real spectrum of higher level of a commutative ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):449–462, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1992:DMF

[167] J. M. Borwein and A. S. Lewis. Decomposition of multivariate functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 463–482, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ghoussoub:1992:SST

[168] N. Ghoussoub, B. Maurey, and W. Schachermayer. Slicings, selections and their applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):483–504, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grzaslewicz:1992:EEP

[169] R. Grzaślewicz, H. Hudzik, and W. Kurc. Extreme and exposed points in Orlicz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):505–515, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1992:PFG

[170] C. K. Gupta, N. D. Gupta, and V. A.

Roman’ckov. Primitivity in free groups and free metabelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):

516–523, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallianpur:1992:SLT

[171] Gopinath Kallianpur and Itaru Mitoma. A Segal–Langevin type stochastic differential equation on a space of generalized functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):524–552, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1992:LIC

[172] Charles Lanski. Lie ideals and central identities with derivation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 553–560, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lubinsky:1992:HHF

[173] D. S. Lubinsky and P. Rabinowitz. Hermite and Hermite–Fejér interpolation and associated product integration rules on the real line: the L_{1} theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 561–590, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laubenbacher:1992:DG

[174] Reinhard C. Laubenbacher and Bruce A. Magurn. SK_{2} and K_{3} of dihedral groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):591–623, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sawyer:1992:HES

[175] P. Sawyer. The heat equation on the spaces of positive definite matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):624–651, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1992:FPT

[176] James C. S. Wong. Fixed point theorems for measurable semigroups of operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):652–664, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1992:SSA

[177] Chang Chang Xi. The structure of Schur algebras S_{k}(n, p) for n ≥ p. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 665–672, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aldaz:1992:CUL

[178] J. M. Aldaz. A characterization of universal Loeb measurability for completely regular Hausdorff spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 673–690, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexopoulos:1992:AHT

[179] G. Alexopoulos. An application of homogenization theory to harmonic analysis: Harnack inequalities and Riesz transforms on Lie groups of polynomial growth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):691–727, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ancona:1992:DFR

[180] Vincenzo Ancona and Bernard Gaveau. Differential forms and resolutions on certain analytic spaces. II. flat resolutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):728–749, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Badertscher:1992:CHP

[181] Erich Badertscher and Tom H. Koornwinder. Continuous Hahn polynomials of differential operator argument and analysis on Riemannian symmetric spaces of constant curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 750–773, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burton:1992:PDE

[182] R. M. Burton and A. R. Dabrowski. Positive dependence of exchangeable sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):774–783, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ding:1992:SVG

[183] Xie Ping Ding and Kok-Keong Tan. A set-valued generalization of Fan’s best approximation theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 784–796, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorkin:1992:BAD

[184] Pamela Gorkin, Keiji Izuchi, and Raymond Mortini. Bourgain algebras of Douglas algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):797–804, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izuchi:1992:ASP

[185] Keiji Izuchi. Analysis on sparse parts in the maximal ideal space of H^{∞}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 805–823, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Louboutin:1992:CNR

[186] S. Louboutin, R. A. Mollin, and H. C.

Williams. Class numbers of real quadratic fields, continued fractions, reduced ideals, prime-producing quadratic polynomials and quadratic residue covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):824–842, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mekler:1992:DCA

[187] Alan Mekler, Matatyahu Rubin, and Charles Steinhorn. Dedekind completeness and the algebraic complexity of o-minimal structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):843–855, ????

1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Srivastav:1992:EPP

[188] Anand Srivastav. Extreme points of positive functionals and spectral states on real Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):856–866, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Swarttouw:1992:ATS

[189] René F. Swarttouw. An addition theorem and some product formulas for the Hahn–Exton q-Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 867–879, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Takahashi:1992:FPT

[190] Wataru Takahashi. Fixed point theorem and nonlinear ergodic theorem for nonexpansive semigroups without convexity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):880–887, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Violette:1992:FPP

[191] Donald Violette and Gilles Fournier. Fixed point principles for cones of a Banach space for the multivalued maps differentiable at the origin and infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):888–896, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexopoulos:1992:LEC

[192] G. Alexopoulos. A lower estimate for central probabilities on polycyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):897–910, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andler:1992:BOT

[193] Martin Andler. Bruhat order and transfer for complex reductive groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 911–923, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ditzian:1992:AFP

[194] Z. Ditzian and D. Jiang. Approximation of functions by polynomials in C[−1, 1]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):924–940, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1992:APS

[195] P. M. Gauthier. Approximation by (pluri)subharmonic functions: fusion and localization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):941–950, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodaire:1992:SML

[196] Edgar G. Goodaire. Six Moufang loops of units. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):951–973, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1992:DRW

[197] Benedict H. Gross and Dipendra Prasad. On the decomposition of a representation of SO_{n} when restricted to SO_{n−1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):974–1002, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaczynski:1992:TTT

[198] Tomasz Kaczynski and Jian Hong Wu. A topological transversality theorem for multi-valued maps in locally convex spaces with applications to neutral equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1003–1013, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noussair:1992:PFE

[199] Ezzat S. Noussair, Charles A. Swanson, and Yang Jianfu. Positive finite energy solutions of critical semilinear elliptic problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1014–1029, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rees:1992:FBS

[200] Rolf S. Rees and Douglas R. Stinson. Frames with block size four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1030–1049, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rodl:1992:RPF

[201] V. Rödl and N. Sauer. The Ramsey property for families of graphs which exclude a given graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1050–1060, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Setti:1992:GEH

[202] Alberto G. Setti. Gaussian estimates for the heat kernel of the weighted Laplacian and fractal measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1061–1078, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stembridge:1992:SCI

[203] J. R. Stembridge. Some conjectures for immanants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1079–1099, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tam:1992:TTG

[204] Laying Tam. A Tauberian theorem for the general Euler–Borel summability method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1100–1120, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeKoninck:1992:NGP

[205] J.-M. De Koninck, I. Kátai, and A. Mercier. On the normal growth of prime factors of integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1121–1154, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bhatia:1992:SVS

[206] R. Bhatia and L. Elsner. Symmetries and variation of spectra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1155–1166, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farsi:1992:QA

[207] Carla Farsi and Neil Watling. Quartic algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1167–1191, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferland:1992:CFA

[208] René Ferland, Xavier Fernique, and Gaston Giroux. Compactness of the fluctuations associated with some generalized nonlinear Boltzmann equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1192–1205, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gu:1992:TWN

[209] Dangsheng Gu. Two-weight norm inequality and Carleson measure in weighted Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1206– 1219, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jacquet:1992:RKI

[210] Hervé Jacquet. Relative Kloosterman integrals for GL(3). II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):

1220–1240, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lupton:1992:RHT

[211] Gregory Lupton and Ronald Umble. Rational homotopy types with the rational cohomology algebra of stunted complex projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1241–1261, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1992:PBC

[212] Murray Marshall. The Pierce–Birkhoff conjecture for curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1262–1271, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohar:1992:CLP

[213] Bojan Mohar. Combinatorial local planarity and the width of graph embeddings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1272–1288, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1992:BPA

[214] K. Pan and E. B. Saff. Best polynomial approximation with linear constraints. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44 (??):1289–1302, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pfeffer:1992:MDG

[215] Washek F. Pfeffer and Brian S. Thomson. Measures defined by gages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??): 1303–1316, ???? 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yan:1992:CGH

[216] Zhimin Yan. A class of generalized hypergeometric functions in several variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(6):1317–1338, December 1992. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1993:MGB

[217] G. F. D. Duff. In memoriam: Gilbert de Beauregard Robinson (1906–1992). Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):3–5, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chabert:1993:PVE

[218] Jean-Luc Chabert and Gilbert Gerboud. Polynômes à valeurs entières et binômes de Fermat. (French) [Polynomials with integer values and Fermat binomials]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):6–21, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DiVincenzo:1993:HPC

[219] Onofrio Mario Di Vincenzo and Rosa Sagona. Homogeneous polynomials, centralizers and derivations in rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 22–32, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedman:1993:CAF

[220] Yaakov Friedman and Bernard Russo.

Classification of atomic facially symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):33–87, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1993:NLP

[221] Arjeh M. Cohen and Robert L. Griess, Jr. Nonlocal Lie primitive subgroups of Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):88–103, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinig:1993:WHI

[222] H. P. Heinig and V. D. Stepanov.

Weighted Hardy inequalities for increasing functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):104–116, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ohyama:1993:MCN

[223] Yoshiyuki Ohyama. On the minimal crossing number and the braid index of links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):117–131, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Philos:1993:OHO

[224] Ch. G. Philos, I. K. Purnaras, and Y. G. Sficas. Oscillations in higher-order neutral differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 132–158, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shao:1993:LIL

[225] Qi Man Shao. On the law of the iterated logarithm for infinite-dimensional Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 159–175, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Waldschmidt:1993:MCL

[226] Michel Waldschmidt. Minorations de combinaisons linéaires de logarithmes de nombres algébriques. (French) [Lower bounds of linear combinations of logarithms of algebraic numbers]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 176–224, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brunner:1993:APE

[227] A. M. Brunner, R. G. Burns, and Sheila Oates-Williams. On almost primitive elements of free groups with an application to Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):225–254, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cochran:1993:RDF

[228] W. George Cochran, Joel H. Shapiro, and David C. Ullrich. Random Dirichlet functions: multipliers and smoothness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):255–268, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Csorgo:1993:RMC

[229] M. Csörgö, Z.-Y. Lin, and Q.-M.

Shao. Randomization moduli of continuity for l^{2}-norm squared Ornstein– Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):269–283, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

El-Zahar:1993:DHD

[230] M. El-Zahar and N. W. Sauer. On the divisibility of homogeneous directed graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):284–294, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1993:PMR

[231] Peter Hoffman. Projective and multigraded representations of monomial and multisigned groups. I. graded representations of a twisted product. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 295–339, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liem:1993:AKC

[232] Vo Thanh Liem and Gerard A. Venema.

On the asphericity of knot complements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):340–356, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lim:1993:LTS

[233] Ming Huat Lim. Linear transformations on symmetric spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):357–368, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magurn:1993:BGS

[234] Bruce A. Magurn. Birelative K_{2} of groups of square-free order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 369–379, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nikiel:1993:HIC

[235] J. Nikiel and E. D. Tymchatyn. On homogeneous images of compact ordered spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):380–393, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1993:SQC

[236] Mizan Rahman. Some quadratic and cubic summation formulas for basic hypergeometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):394–411, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmitt:1993:HAC

[237] William R. Schmitt. Hopf algebras of combinatorial structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):

412–428, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:1993:CSE

[238] Qing Zhang. Conditional sequence entropy and mild mixing extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 429–448, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1993:MWA

[239] M. A. Akcoglu and Y. Déniel. Moving weighted averages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):449–469, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asmar:1993:SIM

[240] Nakhlé Asmar, Earl Berkson, and T. A. Gillespie. Spectral integration of Marcinkiewicz multipliers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 470–482, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bresar:1993:MWP

[241] Matej Bresar and Peter Semrl. Mappings which preserve idempotents, local automorphisms, and local derivations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):483–496, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dajani:1993:GRC

[242] Karma Dajani. Generic results for cocycles with values in a semidirect product. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):497–516, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dumir:1993:CGT

[243] V. C. Dumir, R. J. Hans-Gill, and J. B. Wilker. Contributions to a general theory of view-obstruction problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 517–536, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ge:1993:CRM

[244] Zhong Ge. Collapsing Riemannian metrics to Carnot–Carathéodory metrics and Laplacians to sub-Laplacians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):537–553, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hebert:1993:OPI

[245] Michel Hébert. Sur les opérations partielles implicites et leur relation avec la surjectivité des épimorphismes. (French) [On implicit partial operations and their relationship with the surjectivity of epimorphisms]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):554–575, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kato:1993:CWE

[246] Hisao Kato. Continuum-wise expansive homeomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):576–598, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lovric:1993:CPB

[247] Miroslav Lovrić. Curvature pinching based on integral norms of the curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):599–611, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McLaughlin:1993:SOG

[248] D. McLaughlin, R. Poliquin, J. Vanderwerff, and V. Zizler. Second-order Gateaux differentiable bump functions and approximations in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):612–625, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ninomiya:1993:CSS

[249] Yasushi Ninomiya. Correction to:

“Structure of p-solvable groups with three p-regular classes”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):626, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [95].

Olivier:1993:PPZ

[250] P. F. Olivier and A. O. Watt. Polynomials with a prescribed zero and the Bernstein’s inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):627–637, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahe:1993:CGB

[251] M. Rahe. On a class of generalized baker’s transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 638–649, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sankaran:1993:GAF

[252] P. Sankaran and K. Varadarajan. Group actions on flag manifolds and cobordism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):650–661, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zelewski:1993:HET

[253] Piotr M. Zelewski. On the Hermitian– Einstein tensor of a complex homogeneous vector bundle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):662–672, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschhorn:1993:CAJ

[254] Michael Hirschhorn, Frank Garvan, and Jon Borwein. Cubic analogues of the Jacobian theta function θ(z, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 673–694, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bresar:1993:CPM

[255] Matej Bresar and C. Robert Miers.

Commutativity preserving mappings of von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):695–708, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edward:1993:JMS

[256] Julian Edward. Jacobi matrices and the spectrum of the Neumann operator on a family of Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 709–726, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farkas:1993:STG

[257] Daniel R. Farkas, C. D. Feustel, and Edward L. Green. Synergy in the theories of Gröbner bases and path algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):727–739, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1993:DCU

[258] P. M. Gauthier and N. N. Tarkhanov.

Degenerate cases of uniform approximation by solutions of systems with surjective symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):740–757, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gimigliano:1993:IVS

[259] A. Gimigliano and A. Lorenzini. On the ideal of Veronesean surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 758–777, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herz:1993:RLG

[260] Carl Herz. Representations of Lie groups by contact transformations. II. noncompact simple groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):778–802, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kechagias:1993:MOM

[261] Nondas E. Kechagias. Mod odd modular coinvariants, homology operations, and limit spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):803–819, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liu:1993:QMT

[262] Ke Qin Liu. Quantizations of the module of tensor fields over the Witt algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):820–834, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liu:1993:PSS

[263] Lin Qi Liu. Propagation of singularities for semilinear hyperbolic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):835–846, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mai:1993:ARF

[264] Liem Mai. The analytic rank of a family of elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):847–862, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Okada:1993:ABS

[265] Tatsuya Okada. Asymptotic behavior of skew conditional heat kernels on graph networks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):863–878, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poliquin:1993:CED

[266] R. A. Poliquin and R. T. Rockafellar. A calculus of epi-derivatives applicable to optimization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):879–896, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bartoszynski:1993:CGB

[267] Tomek Bartoszyński, Winfried Just, and Marion Scheepers. Covering games and the Banach–Mazur game: K-tactics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):897–929, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Godoy:1993:OPR

[268] E. Godoy and F. Marcellán. Orthogonal polynomials and rational modifications of measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):930–943, ????

1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goerss:1993:PIH

[269] Paul G. Goerss. Projective and injective Hopf algebras over the Dyer– Lashof algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):944–976, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hager:1993:FDA

[270] Anthony W. Hager and Jorge Martinez. Fraction-dense algebras and spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):977–996, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petechuk:1993:IBL

[271] Vasilij M. Petechuk. Isomorphisms between linear groups over division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (??):997–1008, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peters:1993:ATA

[272] J. R. Peters, Y. T. Poon, and B. H. Wagner. Analytic TAF algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1009–1031, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [311].

Raeburn:1993:DDC

[273] Iain Raeburn and Dana P. Williams. Dixmier–Douady classes of dynamical systems and crossed products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1032–1066, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderPoorten:1993:SCF

[274] A. J. van der Poorten and J. Shallit.

A specialised continued fraction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1067–1079, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stroethoff:1993:ECT

[275] Karel Stroethoff. Essentially commuting Toeplitz operators with harmonic symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1080–1093, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1993:OCS

[276] James S. W. Wong. Oscillation criteria for second order nonlinear differential equations involving integral averages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1094–1103, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yezza:1993:SCS

[277] A. Yezza. Sensitivity and controllability of systems governed by integral equations via proximal analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1104–1120, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1993:CFH

[278] J. M. Borwein and M. Fabián. On convex functions having points of Gâteaux differentiability which are not points of Fréchet differentiability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1121– 1134, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bryant:1993:DSA

[279] Roger M. Bryant and Vesselin Drensky. Dense subgroups of the automorphism groups of free algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1135–1154, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carro:1993:BSI

[280] María J. Carro and Javier Soria. Boundedness of some integral operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1155–1166, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clarke:1993:SCM

[281] F. H. Clarke, R. J. Stern, and P. R. Wolenski. Subgradient criteria for monotonicity, the Lipschitz condition, and convexity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(6):1167–1183, December 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cordes:1993:AWR

[282] Craig M. Cordes. Abstract Witt rings when certain binary forms represent exactly four elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1184–1199, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1993:EFK

[283] Benjamin L. Fine and Georgia Triantafillou. On the equivariant formality of Kähler manifolds with finite group action. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1200–1210, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mao:1993:RKI

[284] Zhengyu Mao. Relative Kloosterman integrals for GL(3). III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1211–1230, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin-Reyes:1993:C

[285] F. J. Martín-Reyes, L. Pick, and A. de la Torre. A_{∞}^{+} condition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(6):1231– 1244, December 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mori:1993:LDL

[286] Tamás F. Móri. The a.s. limit distribution of the longest head run. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):

1245–1262, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mueller:1993:MDW

[287] C. Mueller. A modulus for the 3dimensional wave equation with noise: dealing with a singular kernel. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??): 1263–1275, ???? 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parshall:1993:CQG

[288] Brian Parshall and Jian Pan Wang. Cohomology of quantum groups: the quantum dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1276–1298, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prabaharan:1993:EIW

[289] K. Prabaharan. Existence of invariant weak units in Banach lattices: countably generated left amenable semigroup of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1299–1312, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reid:1993:DMR

[290] Les Reid and Xingping Sun. Distance matrices and ridge function interpolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):1313–1323, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmuland:1993:NOU

[291] Byron Schmuland. Nonsymmetric Ornstein–Uhlenbeck processes in Banach space via Dirichlet forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(6):1324– 1338, December 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1994:HDL

[292] G. Grätzer and H. Lakser. Homomorphisms of distributive lattices as restrictions of congruences. II. planarity and automorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):3–54, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brenti:1994:LZC

[293] Francesco Brenti, Gordon F. Royle, and David G. Wagner. Location of zeros of chromatic and related polynomials of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):55–80, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Csorgo:1994:KGT

[294] Miklós Csörgö, Zheng Yan Lin, and Qi Man Shao. Kernel generated twotime parameter Gaussian processes and some of their path properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 81–119, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Evans:1994:MVB

[295] Steven N. Evans and Edwin A. Perkins. Measure-valued branching diffusions with singular interactions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):120–168, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurrelbrink:1994:CGR

[296] Jürgen Hurrelbrink. Circulant graphs and 4-ranks of ideal class groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 169–183, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenblatt:1994:NCE

[297] Joseph Rosenblatt. Norm convergence in ergodic theory and the behavior of Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):184–199, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanderwerff:1994:MBC

[298] J. Vanderwerff, J. H. M. Whitfield, and V. Zizler. Markusevic bases and Corson compacta in duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):200–211, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luo:1994:NSC

[299] Xue Bo Luo. Necessary and sufficient conditions for hypoellipticity for a class of convolution operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):

212–224, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1994:DPS

[300] Stephen D. Cohen. Dickson polynomials of the second kind that are permutations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):225–238, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Colbourn:1994:SOS

[301] Charles J. Colbourn, Peter B. Gibbons, Rudolf Mathon, Ronald C. Mullin, and Alexander Rosa. The spectrum of orthogonal Steiner triple systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 239–252, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Didierjean:1994:EHC

[302] G. Didierjean and A. Legrand. Équivalences d’homotopie et crochet de Whitehead. (French) [Equivalence of Whitehead’s homotopy and hook]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):253–273, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dowling:1994:GSV

[303] Patrick N. Dowling, Zhibao Hu, and Mark A. Smith. Geometry of spaces of vector-valued harmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):274–283, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erbe:1994:ONP

[304] L. H. Erbe and Qingkai Kong. Oscillation and nonoscillation properties of neutral differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 284–297, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1994:SAA

[305] C. K. Gupta, N. D. Gupta, and G. A. Noskov. Some applications of Artamonov–Quillen–Suslin theorems to metabelian inner rank and primitivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):298–307, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herb:1994:UOI

[306] Rebecca A. Herb. Unipotent orbital integrals of Hecke functions for GL(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):308–323, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holland:1994:CTI

[307] D. Holland. Chinburg’s third invariant in the factorisability defect class group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):324–342, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1994:MWE

[308] Roger L. Jones and James Olsen. Multiparameter weighted ergodic theorems.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):343–356, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreuzer:1994:CMD

[309] Martin Kreuzer. On the canonical module of a 0-dimensional scheme. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 357–379, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parameswaran:1994:NTT

[310] Mangalam R. Parameswaran. New Tauberian theorems from old. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 380–394, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peters:1994:CAT

[311] J. R. Peters, Y. T. Poon, and B. H.

Wagner. Correction to: “Analytic TAF algebras”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):395–396, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [272].

Poon:1994:OAC

[312] Yiu Tung Poon and Zhong-Jin Ruan.

Operator algebras with contractive approximate identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):397–414, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shiga:1994:TCP

[313] Tokuzo Shiga. Two contrasting properties of solutions for one-dimensional stochastic partial differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):415–437, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:1994:DLGa

[314] N. Th. Varopoulos. Diffusion on Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):438–448, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Acquistapace:1994:WDS

[315] F. Acquistapace, F. Broglia, and E. Fortuna. When is a distribution of signs locally completable? Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):449–473, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aulaskari:1994:AAF

[316] Rauno Aulaskari and Peter Lappan. Additive automorphic functions and Bloch functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):474–484, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Botvinnik:1994:SHT

[317] Boris I. Botvinnik and Stanley O. Kochman. Singularities and higher torsion in symplectic cobordism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):

485–516, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1994:NFF

[318] Paul M. Cohn and Christophe Reutenauer.

A normal form in free fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 517–531, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farid:1994:ISO

[319] F. O. Farid and K. Varadarajan. Isometric shift operators on C(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 532–542, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huang:1994:RPR

[320] John Q. Huang. Real projective representations of S_{N} and A_{N}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 543–573, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Komorowski:1994:ICS

[321] Ryszard Komorowski. Isometric characterizations of l_{p}^{n} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):574–585, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lakey:1994:WFT

[322] Joseph D. Lakey. Weighted Fourier transform inequalities via mixed norm Hausdorff–Young inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 586–601, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lancaster:1994:INS

[323] P. Lancaster and L. Rodman. Invariant neutral subspaces for symmetric and skew real matrix pairs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):602–618, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leviatan:1994:DAR

[324] D. Leviatan and D. S. Lubinsky. Degree of approximation by rational functions with prescribed numerator degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):619–633, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Page:1994:DSI

[325] Stanley S. Page and Yi Qiang Zhou. On direct sums of injective modules and chain conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):634–647, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putcha:1994:CCA

[326] Mohan S. Putcha. Conjugacy classes in algebraic monoids. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):648–661, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Young:1994:LPM

[327] Wo-Sang Young. Littlewood–Paley and multiplier theorems for Vilenkin– Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):662–672, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arkowitz:1994:CSM

[328] Martin Arkowitz and Marek Golasiński.

Co-H-structures on Moore spaces of type (G, 2). Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):673–686, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1994:EIN

[329] Peter B. Borwein. Exact inequalities for the norms of factors of polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):687–698, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokovic:1994:CCM

[330] Dragomir Ž. – Doković and Nguyêñ Quôć Thǎńg. Conjugacy classes of maximal tori in simple real algebraic groups and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(4):699–717, August 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [359].

Drensky:1994:CCH

[331] Vesselin Drensky and Antonio Giambruno. Cocharacters, codimensions and Hilbert series of the polynomial identities for 2 × 2 matrices with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):718–733, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garrido:1994:HFA

[332] M. I. Garrido, J. Gómez Gil, and J. A. Jaramillo. Homomorphisms on function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):734–745, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hilgert:1994:CTB

[333] Joachim Hilgert. A convexity theorem for boundaries of ordered symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):746–757, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hindman:1994:SPM

[334] Neil Hindman, Jimmie Lawson, and Amha Lisan. Separating points of βN by minimal flows. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(4):758–771, August 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:1994:FBP

[335] Zaiqing Li. Formulas for Brown– Peterson operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):772–792, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McClanahan:1994:SRA

[336] Kevin McClanahan. Simplicity of reduced amalgamated products of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):793–807, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miao:1994:EIM

[337] Tianxuan Miao. The extensions of an invariant mean and the set LIMsimTLIM. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):808–817, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:1994:RER

[338] N. Christopher Phillips. Reduction of exponential rank in direct limits of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):818–853, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Quinn:1994:PDP

[339] Terrance Quinn. Products of decomposable positive operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):854–871, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruttimann:1994:WPF

[340] Gottfried T. Rüttimann. Weakly purely finitely additive measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 872–885, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1994:ESG

[341] Tara L. Smith. Extra-special groups of order 32 as Galois groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 886–896, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akhiezer:1994:CHS

[342] D. N. Akhiezer and B. Gilligan. On complex homogeneous spaces with top homology in codimension two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 897–919, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabbri:1994:RQH

[343] Marc A. Fabbri and Frank Okoh. Representations of quantum Heisenberg algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):920–929, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1994:IR

[344] Benedict H. Gross and Dipendra Prasad. On irreducible representations of SO_{2n+1} × SO_{2m}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):930–950, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haddad:1994:FCC

[345] L. Haddad and I. G. Rosenberg. Finite clones containing all permutations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):951–970, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jain:1994:WPW

[346] S. K. Jain, S. R. López-Permouth, K. Oshiro, and M. A. Saleh. Weakly projective and weakly injective modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):971–981, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kiesel:1994:TCT

[347] Rüdiger Kiesel and Ulrich Stadtmüller.

Tauberian- and convexity theorems for certain (N, p, q)-means. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):982–994, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lari-Lavassani:1994:STS

[348] Ali Lari-Lavassani and Yung-Chen Lu.

The stability theorems for subgroups of scrA and scrK. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):995–1006, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morenz:1994:SCS

[349] Phillip B. Morenz. The structure of C^{∗}-convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1007–1026, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagy:1994:LIS

[350] Péter T. Nagy and Karl Strambach.

Loops as invariant sections in groups, and their geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1027–1056, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salvador:1994:WLN

[351] P. Ortega Salvador. Weighted Lorentz norm inequalities for the one-sided Hardy–Littlewood maximal functions and for the maximal ergodic operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):1057–1072, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:1994:DLGb

[352] N. Th. Varopoulos. Diffusion on Lie groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1073–1092, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wehrung:1994:UTO

[353] Friedrich Wehrung. The universal theory of ordered equidecomposability types semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1093–1120, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abdulali:1994:ACF

[354] Salman Abdulali. Algebraic cycles in families of abelian varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1121–1134, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Albert:1994:UVG

[355] Michael H. Albert and John Lawrence.

Unification in varieties of groups: nilpotent varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1135–1149, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1994:QME

[356] Lawrence G. Brown, James A. Mingo, and Nien-Tsu Shen. Quasi-multipliers and embeddings of Hilbert C^{∗}-bimodules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46 (??):1150–1174, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bula:1994:KBP

[357] W. Bula, J. Nikiel, and E. D. Tymchatyn. The Königsberg Bridge Problem for Peano continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1175–1187, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ciesielski:1994:TDC

[358] Krzysztof Ciesielski. Topologizing different classes of real functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1188–1207, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokovic:1994:CCC

[359] D. Ž. – Doković and Nguyêñ Q. Thǎńg.

Correction to: “Conjugacy classes of maximal tori in simple real algebraic groups and applications”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1208– 1210, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [330].

Farsi:1994:ACF

[360] Carla Farsi and Neil Watling. Abstract characterizations of fixed point subalgebras of the rotation algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1211–1237, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iscoe:1994:AED

[361] I. Iscoe, D. McDonald, and K. Qian. Asymptotics of the exit distribution for Markov jump processes; application to ATM. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1238–1262, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jaworski:1994:EBA

[362] Wojciech Jaworski. Exponential boundedness and amenability of open subsemigroups of locally compact groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1263–1274, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

May:1994:LBR

[363] Coy L. May. A lower bound for the real genus of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1275–1286, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nielsen:1994:SYI

[364] Ole A. Nielsen. Sharpness in Young’s inequality for convolution products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??): 1287–1298, ???? 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tokunaga:1994:DGC

[365] Hiro o Tokunaga. On dihedral Galois coverings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):1299–1317, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1994:AMR

[366] R. Wong and J.-M. Zhang. Asymptotic monotonicity of the relative extrema of Jacobi polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(6):1318– 1337, December 1994. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baillon:1995:AON

[367] J.-B. Baillon, Ph. Clément, A. Greven, and F. den Hollander. On the attracting orbit of a non-linear transformation arising from renormalization of hierarchically interacting diffusions. I. the compact case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):3–27, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berenstein:1995:GCI

[368] Carlos A. Berenstein, Bao Qin Li, and Alekos Vidras. Geometric characterization of interpolating varieties for the (FN)-space A_{p}^{0} of entire functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 28–43, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blasco:1995:MSA

[369] Oscar Blasco. Multipliers on spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):44–64, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1995:ZII

[370] Peter B. Borwein, Weiyu Chen, and Karl Dilcher. Zeros of iterated integrals of polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):65–87, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duran:1995:OPR

[371] Antonio J. Duran. On orthogonal polynomials with respect to a positive definite matrix of measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):88–112, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoechsmann:1995:UIG

[372] Klaus Hoechsmann. Units in integral group rings for order pq. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):113–131, ????

1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koppelberg:1995:DUB

[373] Sabine Koppelberg and Saharon Shelah. Densities of ultraproducts of Boolean algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):132–145, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Riley:1995:RLA

[374] D. M. Riley and A. Shalev. Restricted Lie algebras and their envelopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 146–164, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rockner:1995:QRD

[375] Michael Röckner and Byron Schmuland.

Quasi-regular Dirichlet forms: examples and counterexamples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):165–200, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sakuma:1995:HAC

[376] Makoto Sakuma. Homology of abelian coverings of links and spatial graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):201–224, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asmar:1995:CEG

[377] Nakhlé Asmar, Earl Berkson, and T. A.

Gillespie. Convolution estimates and generalized de Leeuw theorems for multipliers of weak type (1, 1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 225–245, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beidleman:1995:GWS

[378] James C. Beidleman, Martyn R. Dixon, and Derek J. S. Robinson. The generalized Wielandt subgroup of a group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):246–261, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1995:ATB

[379] Jonathan M. Borwein and Roland Girgensohn. Addition theorems and binary expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):262–273, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bovdi:1995:TGR

[380] Victor Bovdi. Twisted group rings whose units form an FC-group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 274–289, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clauss:1995:ECI

[381] Jon M. Clauss. Elementary chains of invariant subspaces of a Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):290–301, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Darmon:1995:TMC

[382] Henri Darmon. Thaine’s method for circular units and a conjecture of Gross. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):302–317, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Debs:1995:TPS

[383] Gabriel Debs, Gilles Godefroy, and Jean Saint-Raymond. Topological properties of the set of norm-attaining linear functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):318–329, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fernandez-Cabrera:1995:EAE

[384] L. M. Fernández-Cabrera, F. J. Martín-Reyes, and J. L. Torrea. On the ergodic averages and the ergodic Hilbert transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):330–343, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldberg:1995:RNC

[385] David Goldberg. Reducibility for nonconnected p-adic groups, with G^{∘} of prime index. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):344–363, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hamel:1995:LPS

[386] A. M. Hamel and I. P. Goulden. Lattice paths and a Sergeev–Pragacz formula for skew supersymmetric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 364–382, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herfort:1995:FPA

[387] Wolfgang N. Herfort, Luis Ribes, and Pavel A. Zalesskii. Fixed points of automorphisms of free pro-p groups of rank 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):383–404, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huang:1995:BRP

[388] Xiao Jun Huang. A boundary rigidity problem for holomorphic mappings on some weakly pseudoconvex domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):405–420, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huang:1995:FPI

[389] Houben Huang and Timothy D. Hudson. Factorization of positive invertible operators in AF algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):421–435, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koelink:1995:AFB

[390] H. T. Koelink. Addition formula for big q-legendre polynomials from the quantum SU(2) group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):436–448, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binding:1995:DNS

[391] Paul Binding, D. R. Farenick, and Chi-Kwong Li. A dilation and norm in several variable operator theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 449–461, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cheng:1995:ARF

[392] R. Cheng. Analytic range functions of several variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):462–473, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deMedicis:1995:GSN

[393] Anne de Médicis and Pierre Leroux.

Generalized Stirling numbers, convolution formulae and p, q-analogues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 474–499, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haddock:1995:CTL

[394] John R. Haddock, Shi Gui Ruan, Jian Hong Wu, and Hua Xing Xia. Comparison theorems of Liapunov– Razumikhin type for NFDEs with infinite delay. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):500–526, ????

1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hwang:1995:BGA

[395] Yoon Sung Hwang and Bill Jacob. Brauer group analogues of results relating the Witt ring to valuations and Galois theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):527–543, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langevin:1995:CHS

[396] Rémi Langevin, Gilbert Levitt, and Harold Rosenberg. Classes d’homotopie de surfaces avec rebroussements et queues d’aronde dans R^{3}. (French) [Homotopy classes of surfaces with cusps and dovetails in R^{3}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):544–572, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moody:1995:VDD

[397] R. V. Moody and J. Patera. Voronoi domains and dual cells in the generalized kaleidoscope with applications to root and weight lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):573–605, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murnaghan:1995:LCE

[398] Fiona Murnaghan. Local character expansions for supercuspidal representations of U(3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):606–640, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schulte:1995:FEC

[399] Egon Schulte and Asia Ivić Weiss. Free extensions of chiral polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 641–654, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shen:1995:FRT

[400] Yuan Yuan Shen and Lawrence C. Washington. A family of real p^{n}-tic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):655–672, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aron:1995:WSC

[401] R. M. Aron, B. J. Cole, and T. W.

Gamelin. Weak-star continuous analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):673–683, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fang:1995:ECC

[402] G. Fang. On the existence and the classification of critical points for nonsmooth functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):684–717, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felipe:1995:RCC

[403] Maria Jose Felipe and Gabriel Navarro.

Relative character correspondences in finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):718–727, ????

1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferrando:1995:MET

[404] S. E. Ferrando. Moving ergodic theorems for superadditive processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 728–743, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fowler:1995:FS

[405] Neal J. Fowler. Free E_{0}-semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):744–785, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jamjoom:1995:TPA

[406] Fatmah B. Jamjoom. On the tensor products of JW-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 786–800, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:1995:SAD

[407] Tzee Char Kuo. A simple algorithm for deciding primes in K[![x, y]!]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 801–816, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ma:1995:MPA

[408] Zhi Ming Ma and Michael Röckner. Markov processes associated with positivity preserving coercive forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 817–840, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacDonald:1995:DPN

[409] Gordon W. MacDonald. Distance from projections to nilpotents. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 841–851, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McIntosh:1995:EAW

[410] David I. McIntosh. Ergodic averages for weight functions moved by non-linear transformations on R^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):852–876, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xu:1995:CCV

[411] Yeren Xu. Continuation of complex varieties across rectifiable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 877–896, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berndt:1995:SVR

[412] Bruce C. Berndt and Heng Huat Chan.

Some values for the Rogers–Ramanujan continued fraction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):897–914, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruck:1995:CIC

[413] Rainer Brück and Jürgen Müller. Closed ideals in a convolution algebra of holomorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):915–928, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellers:1995:CIR

[414] Harald Ellers. Cliques of irreducible representations, quotient groups, and Brauer’s theorems on blocks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 929–945, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Girardin:1995:EMG

[415] Valérie Girardin. Extreme measures with given moments or marginals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 946–958, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodman:1995:BST

[416] T. N. T. Goodman and A. Sharma. A Bernstein–Schoenberg type operator: shape preserving and limiting behaviour. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):959–973, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hales:1995:FLS

[417] Thomas C. Hales. On the fundamental lemma for standard endoscopy: reduction to unit elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):974–994, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewis:1995:CMS

[418] A. S. Lewis. Consistency of moment systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):995–1006, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longstaff:1995:PSS

[419] W. E. Longstaff and H. Radjavi. On permutability and submultiplicativity of spectral radius. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1007–1022, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mollin:1995:PDA

[420] R. A. Mollin and H. C. Williams. Proof, disproof and advances concerning certain conjectures on real quadratic fields oldQ(\sqrt{N^{2} + 4}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1023–1036, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Passare:1995:RIT

[421] Mikael Passare and August Tsikh.

Residue integrals and their Mellin transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1037–1050, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ron:1995:FSB

[422] Amos Ron and Zuowei Shen. Frames and stable bases for shift-invariant subspaces of L_{2}(R^{d}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1051–1094, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Welland:1995:FOT

[423] Grant Welland and Shi Ying Zhao. epsilon-families of operators in Triebel– Lizorkin and tent spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1095– 1120, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baratchart:1995:RIE

[424] L. Baratchart, E. B. Saff, and F. Wielonsky. Rational interpolation of the exponential function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1121–1147, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1995:PSS

[425] H. H. Brungs and M. Schröder. Prime segments of skew fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1148– 1176, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cascante:1995:CMS

[426] Carme Cascante and Joaquín M. Ortega. Carleson measures on spaces of Hardy–Sobolev type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1177–1200, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cummins:1995:RHR

[427] C. J. Cummins and S. P. Norton. Rational Hauptmoduls are replicable. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 1201–1218, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gingold:1995:APP

[428] H. Gingold and A. Knopfmacher. Analytic properties of power product expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1219–1239, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Joseph:1995:TNT

[429] James E. Joseph and Myung H. Kwack.

The topological nature of two Noguchi theorems on sequences of holomorphic mappings between complex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):1240–1252, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ledet:1995:GGG

[430] Arne Ledet. On 2-groups as Galois groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1253–1273, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minac:1995:DWR

[431] Ján Minác and Tara L. Smith. Decomposition of Witt rings and Galois groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):1274–1289, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Movahedi-Lankarani:1995:TQ

[432] H. Movahedi-Lankarani and R. Wells. The topology of quasibundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??): 1290–1316, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tie:1995:I

[433] Jingzhi Tie. Imbedding C^{1} into H_{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47 (??):1317–1328, ???? 1995. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:1995:CVV

[434] Jae-Hyun Yang. Construction of vector valued modular forms from Jacobi forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(6):1329–1339, December 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andruskiewitsch:1996:NEH

[435] Nicolás Andruskiewitsch. Notes on extensions of Hopf algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 3–42, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:1996:HPC

[436] Christian Berg and Mourad E. H. Ismail. q-hermite polynomials and classical orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):43–63, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Botvinnik:1996:MPS

[437] Boris Botvinnik and Peter B. Gilkey.

Metrics of positive scalar curvature on spherical space forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):64–80, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fouvry:1996:DSP

[438] Etienne Fouvry and M. Ram Murty. On the distribution of supersingular primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):81–104, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guo:1996:GRW

[439] Jiandong Guo. On a generalization of a result of Waldspurger. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):105–142, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ORegan:1996:SSS

[440] Donal O’Regan. Solvability of some singular boundary value problems on the semi-infinite interval. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):143–158, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Packer:1996:MCC

[441] Judith A. Packer. Moore cohomology and central twisted crossed product C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):159–174, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1996:ASS

[442] Ian F. Putnam. C^{∗}-algebras from Smale spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):175–195, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosen:1996:PPP

[443] Lon Rosen. Positive powers of positive positive-definite matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 196–209, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Voit:1996:CAD

[444] Michael Voit. Compact almost discrete hypergroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):210–224, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aleman:1996:BSD

[445] Alexandru Aleman, Stefan Richter, and William T. Ross. Bergman spaces on disconnected domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):225–243, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bakaev:1996:PES

[446] Nikolai Bakaev. A priori estimates for some classes of difference schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):244–257, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barr:1996:AM

[447] Michael Barr. Acyclic models. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 258–273, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1996:PLS

[448] Liang Chen. Polynomial Lie subalgebras of the infinite matrix Lie algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 274–287, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cima:1996:RAF

[449] Joseph A. Cima and Michael Stessin. On the recovery of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 288–301, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dooley:1996:FGR

[450] A. H. Dooley and S. J. Eigen. A family of generalized Riesz products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 302–315, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Doukhan:1996:SRP

[451] P. Doukhan and F. Gamboa. Superresolution rates in Prokhorov metric. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 316–329, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gootman:1996:SIT

[452] Elliot C. Gootman and Aldo J. Lazar. Spectra for infinite tensor product type actions of compact groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 330–342, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1996:EFL

[453] Donggyun Kim. Explicit forms of local lifting for GL_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):343–362, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lu:1996:HLS

[454] Shanzhen Lu and Dachun Yang. Hardy– Littlewood–Sobolev theorems of fractional integration on Herz-type spaces and its applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):363–380, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olin:1996:SOD

[455] Robert F. Olin and Liming Yang. A subnormal operator and its dual. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):

381–396, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sjostrand:1996:DRS

[456] Johannes Sjöstrand. Density of resonances for strictly convex analytic obstacles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):397–448, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allison:1996:CQC

[457] Bruce N. Allison and Yun Gao. Central quotients and coverings of Steinberg unitary Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):449–482, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arlettaz:1996:EHG

[458] Dominique Arlettaz. The exponent of the homotopy groups of Moore spectra and the stable Hurewicz homomorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):483–495, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bonilla:1996:BPA

[459] A. Bonilla and R. Trujillo-González.

Bounded pointwise approximation of solutions of elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 496–511, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brudern:1996:APP

[460] Jörg Brüdern and Alberto Perelli. The addition of primes and power. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 512–526, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chua:1996:WSS

[461] Seng-Kee Chua. On weighted Sobolev spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):527–541, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fleischmann:1996:STT

[462] Klaus Fleischmann, Jürgen Gärtner, and Ingemar Kaj. A Schilder type theorem for super-Brownian motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 542–568, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:1996:MLO

[463] I. P. Goulden and D. M. Jackson. Maps in locally orientable surfaces, the double coset algebra, and zonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):569–584, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holay:1996:GID

[464] Sandeep H. Holay. Generators of ideals defining certain surfaces in projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):585–595, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1996:RFM

[465] Charles Lanski. Rings with finite maximal invariant subrings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):596–606, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mankiewicz:1996:SPW

[466] Piotr Mankiewicz and Nicole Tomczak-Jaegermann. Structural properties of weak cotype 2 spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):607–624, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Randrianantoanina:1996:CCP

[467] Narcisse Randrianantoanina. Complemented copies of l^{1} and Pelczyński’s property (V^{∗}) in Bochner function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):625–640, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:1996:DLG

[468] N. Th. Varopoulos. Diffusion on Lie groups. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):641–672, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

An:1996:DCC

[469] Jianbei An. Dade’s conjecture for Chevalley groups Gsb2(q) in nondefining characteristics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):673–691, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brenken:1996:LPS

[470] Berndt Brenken. The local product structure of expansive automorphisms of solenoids and their associated C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):692–709, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Damelin:1996:NSCa

[471] S. B. Damelin and D. S. Lubinsky. Necessary and sufficient conditions for mean convergence of Lagrange interpolation for erdős weights. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):710–736, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Damelin:1996:NSCb

[472] S. B. Damelin and D. S. Lubinsky. Necessary and sufficient conditions for mean convergence of Lagrange interpolation for erdős weights. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):737–757, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fegan:1996:FOO

[473] H. D. Fegan and B. Steer. First order operators on manifolds with a group action. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):758–776, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grozev:1996:REF

[474] Georgi R. Grozev and Qazi I. Rahman.

Reconstruction of entire functions from irregularly spaced sample points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 777–793, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hernandez:1996:SRI

[475] Francisco L. Hernandez and Nigel J.

Kalton. Subspaces of rearrangementinvariant spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):794–833, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jourani:1996:ESC

[476] A. Jourani and L. Thibault. Extensions of subdifferential calculus rules in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):834–848, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kossowski:1996:FCC

[477] Marek Kossowski. Fiber completions, contact singularities and single valued solutions for C^{∞}-second order ODE. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):849–870, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:1996:OCS

[478] Horng-Jaan Li and Wei-Ling Liu. Oscillation criteria for second order neutral differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):871–886, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Popescu:1996:PDF

[479] Gelu Popescu. Positive-definite functions on free semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 887–896, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agoston:1996:HDQ

[480] István Ágoston, Vlastimil Dlab, and Erzsébet Lukács. Homological duality and quasi-heredity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):897–917, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1996:FEE

[481] K. Benabdallah and M. A. Ouldbeddi. On finite essential extensions of torsion free abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):918–929, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakazi:1996:RFW

[482] Takahiko Nakazi and Masahiro Yamada. Riesz’s functions in weighted Hardy and Bergman spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):930–945, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1996:TCI

[483] George A. Elliott and Hongbing Su.

K-theoretic classification for inductive limit Z_{2} actions on AF algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 946–958, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldman:1996:PDL

[484] M. L. Gol’dman, H. P. Heinig, and V. D.

Stepanov. On the principle of duality in Lorentz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):959–979, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kishimoto:1996:SSP

[485] Akitaka Kishimoto and Alex Kumjian.

Simple stably projectionless C^{∗}-algebras arising as crossed products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 980–996, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewis:1996:NFR

[486] Mark L. Lewis. Nonabelian fullyramified sections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):997–1017, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delaPena:1996:FSN

[487] J. A. de la Peña and A. Skowroński. Forbidden subcategories of non-polynomial growth tame simply connected algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1018–1043, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ravenel:1996:HR

[488] Douglas C. Ravenel and W. Stephen Wilson. The Hopf ring for P (n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1044–1063, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salim:1996:MGA

[489] Mohamed A. M. Salim and Robert Sandling. The modular group algebra problem for small p-groups of maximal class. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1064–1078, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sellier:1996:AEC

[490] A. Sellier. Asymptotic expansion of a class of multi-dimensional integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1079–1090, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Skowronski:1996:DMN

[491] A. Skowroński and G. Zwara. On degenerations of modules with nondirecting indecomposable summands. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1091–1120, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alzati:1996:NCV

[492] Alberto Alzati, Marina Bertolini, and Gian Mario Besana. Numerical criteria for very ampleness of divisors on projective bundles over an elliptic curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1121–1137, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Basit:1996:NAP

[493] Bolis Basit and Chuanyi Zhang. New almost periodic type functions and solutions of differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1138–1153, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernstein:1996:CM

[494] Daniel J. Bernstein and Jeffrey C. Lagarias. The 3x+1 conjugacy map. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1154–1169, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokuchaev:1996:FSI

[495] Michael A. Dokuchaev and Stanley O.

Juriaans. Finite subgroups in integral group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):1170–1179, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duran:1996:MTO

[496] Antonio J. Duran. Markov’s theorem for orthogonal matrix polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??): 1180–1195, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herman:1996:CBW

[497] Allen Herman. A constructive Brauer–Witt theorem for certain solvable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):1196–1209, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1996:IRP

[498] I. M. Isaacs. Induction and restriction of pi-partial characters and their lifts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):1210–1223, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kapovich:1996:GTQ

[499] Ilya Kapovich and Hamish Short. Greenberg’s theorem for quasiconvex subgroups of word hyperbolic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1224–1244, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1996:RS

[500] Henry H. Kim. The residual spectrum of G_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1245–1272, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miao:1996:CLC

[501] Tianxuan Miao. Compactness of a locally compact group G and geometric properties of A_{p}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):1273–1285, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Russell:1996:SHC

[502] Peter Russell. Sufficiently homogeneous closed embeddings of A^{n−1} into A^{n} are linear. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(??):1286–1295, ???? 1996. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Trnkova:1996:DSC

[503] V. Trnková and J. Sichler. Disciplined spaces and centralizer clone segments. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48 (??):1296–1323, ???? 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Witherspoon:1996:RRT

[504] S. J. Witherspoon. The representation ring of the twisted quantum double of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 48(6):1324– 1338, December 1996. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1997:SPO

[505] Mustafa A. Akcoglu, Dzung M. Ha, and Roger L. Jones. Sweeping out properties of operator sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):3–23, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bertoin:1997:SBP

[506] Jean Bertoin, Jean-François Le Gall, and Yves Le Jan. Spatial branching processes and subordination. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 24–54, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1997:NFE

[507] Huaihui Chen and Paul M. Gauthier. Normal functions: L^{p} estimates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 55–73, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:1997:CAS

[508] Y. K. Hu, K. A. Kopotun, and X. M. Yu. Constrained approximation in Sobolev spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):74–99, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lance:1997:MIS

[509] T. L. Lance and M. I. Stessin. Multiplication invariant subspaces of Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):100–118, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osborn:1997:ALA

[510] J. Marshall Osborn. Automorphisms of the Lie algebras W^{∗} in characteristic 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):119–132, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reeder:1997:EPA

[511] Mark Reeder. Exterior powers of the adjoint representation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):133–159, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rieffel:1997:CLD

[512] Marc A. Rieffel. The classical limit of dynamics for spaces quantized by an action of R^{d}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):160–174, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xu:1997:OPF

[513] Yuan Xu. Orthogonal polynomials for a family of product weight functions on the spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):175–192, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casali:1997:CPM

[514] Maria Rita Casali. Classifying PL 5-manifolds by regular genus: the boundary case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):193–211, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coll:1997:DED

[515] B. Coll, A. Gasull, and R. Prohens. Differential equations defined by the sum of two quasi-homogeneous vector fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):212–231, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edward:1997:STN

[516] Julian Edward. Spectral theory for the Neumann Laplacian on planar domains with horn-like ends. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):232–262, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [697].

Hamel:1997:DFS

[517] A. M. Hamel. Determinantal forms for symplectic and orthogonal Schur functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):263–282, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCall:1997:RCG

[518] Thomas M. McCall, Charles J. Parry, and Ramona R. Ranalli. The 2-rank of the class group of imaginary bicyclic biquadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):283–300, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Merlini:1997:SAC

[519] Donatella Merlini, Douglas G. Rogers, Renzo Sprugnoli, and M. Cecilia Verri.

On some alternative characterizations of Riordan arrays. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):301–320, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mountford:1997:CCT

[520] T. S. Mountford. A complete convergence theorem for attractive reversible nearest particle systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 321–337, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rousseau:1997:LBC

[521] C. Rousseau and B. Toni. Local bifurcations of critical periods in the reduced Kukles system. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):338–358, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sawyer:1997:EHK

[522] P. Sawyer. Estimates for the heat kernel on SL(n, R)/ SO(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):359–372, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stokman:1997:LTB

[523] Jasper V. Stokman and Tom H. Koornwinder. Limit transitions for BC type multivariable orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):373–404, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vulakh:1997:HCF

[524] L. Ya. Vulakh. On Hurwitz constants for Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):405–416, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boe:1997:CCH

[525] Brian D. Boe and Joseph H. G. Fu. Characteristic cycles in Hermitian symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):417–467, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burris:1997:FSLa

[526] Stanley Burris and András Sárközy. Fine spectra and limit laws I. first-order laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):468–498, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fitzgerald:1997:GWR

[527] Robert W. Fitzgerald. Gorenstein Witt rings II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):499–519, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1997:COP

[528] Mourad E. H. Ismail and Dennis Stanton. Classical orthogonal polynomials as moments. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):520–542, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1997:SST

[529] Mourad E. H. Ismail, Mizan Rahman, and Sergei K. Suslov. Some summation theorems and transformations for q-series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):543–567, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mateu:1997:CAH

[530] Joan Mateu. A counterexample in L^{p} approximation by harmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):568–582, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pal:1997:SDQ

[531] Janos Pal and Dana Schlomiuk. Summing up the dynamics of quadratic Hamiltonian systems with a center. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):583–599, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosset:1997:STS

[532] Shmuel Rosset and Alon Wasserman.

The Schreier technique for subalgebras of a free Lie algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):600–616, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stahl:1997:ZSG

[533] Saul Stahl. On the zeros of some genus polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):617–640, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [700].

Burris:1997:FSLb

[534] Stanley Burris, Kevin Compton, Andrew Odlyzko, and Bruce Richmond. Fine spectra and limit laws. II. firstorder 0–1 laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):641–652, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cascante:1997:CMS

[535] Carme Cascante and Joaquin M. Ortega. On q-Carleson measures for spaces of M-harmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):653–674, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deCataldo:1997:SAT

[536] Mark Andrea A. de Cataldo. Some adjunction-theoretic properties of codimension two non-singular subvarities of quadrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):675–695, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Charitos:1997:GFI

[537] Charalambos Charitos and Georgios Tsapogas. Geodesic flow on ideal polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):696–707, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duran:1997:DQT

[538] Antonio J. Duran and Pedro Lopez-Rodriguez. Density questions for the truncated matrix moment problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):708–721, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elder:1997:GMS

[539] G. Griffith Elder and Manohar L.

Madan. Galois module structure of the integers in wildly ramified C_{p}×C_{p} extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):722–735, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fendler:1997:DOP

[540] Gero Fendler. Dilations of one parameter semigroups of positive contractions on L^{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(4):736–748, August 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howe:1997:THW

[541] Lawrence Howe. Twisted Hasse–Weil L-functions and the rank of Mordell– Weil groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):749–771, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jie:1997:FDR

[542] Xiao Jie. Finite dimensional representations of U_{t}(rmsl(2)) at roots of unity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (4):772–787, August 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lichtman:1997:TFR

[543] A. I. Lichtman. Trace functions in the ring of fractions of polycyclic group rings, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):788–797, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yu:1997:BSP

[544] Minqi Yu and Xiting Lian. Boundedness of solutions of parabolic equations with anisotropic growth conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 798–809, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1997:ZDO

[545] K. Pan. The zero distribution of orthogonal rational functions on the unit circle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):810–819, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robart:1997:ISA

[546] Thierry Robart. Sur l’intégrabilité des sous-algèbres de Lie en dimension infinie. (French) [On the integrability of Lie subalgebras of infinite dimension]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):820–839, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rodman:1997:NHS

[547] Leiba Rodman. Non-Hermitian solutions of algebraic Riccati equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):840–854, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1997:RCS

[548] Samuel Bruce Smith. Rational classification of simple function space components for flag manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):855–864, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:1997:MLO

[549] I. P. Goulden and D. M. Jackson. Maps in locally orientable surfaces and integrals over real symmetric surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):865–882, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Okounkov:1997:PCG

[550] Andrei Okounkov. Proof of a conjecture of Goulden and Jackson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 883–886, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1997:PCR

[551] Peter Borwein and Christopher Pinner. Polynomials with {0, +1, −1} coefficients and a root close to a given point. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):887–915, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brylinski:1997:QDN

[552] Ranee Brylinski. Quantization of the 4dimensional nilpotent orbit of SL(3, R). Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):916–943, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jia:1997:AMR

[553] R. Q. Jia, S. D. Riemenschneider, and D. X. Zhou. Approximation by multiple refinable functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):944–962, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:1997:HA

[554] Huaxin Lin. Homomorphisms from C(X) into C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):963– 1009, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorente:1997:CTW

[555] Maria Lorente. A characterization of two weight norm inequalities for onesided operators of fractional type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1010–1033, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saff:1997:RSB

[556] E. B. Saff and H. Stahl. Ray sequences of best rational approximants for |x|^{α}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):1034–1065, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shibata:1997:MVE

[557] Tetsutaro Shibata. Multiparameter variational eigenvalue problems with indefinite nonlinearity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1066–1088, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burke:1997:SWM

[558] Maxim R. Burke and Krzysztof Ciesielski. Sets on which measurable functions are determined by their range. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1089–1116, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:1997:NAL

[559] Zhiguo Hu. The von Neumann algebra VN(G) of a locally compact group and quotients of its subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1117–1138, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kraus:1997:MEP

[560] Alain Kraus. Majorations effectives pour l’équation de Fermat généralisée. (French) [Effective maximizations for the generalized Fermat equation]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1139–1161, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ku:1997:IIS

[561] Hsu-Tung Ku, Mei-Chin Ku, and Xin-Min Zhang. Isoperimetric inequalities on surfaces of constant curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1162–1187, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leen:1997:FIG

[562] Michael J. Leen. Factorization in the invertible group of a C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1188–1205, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Letzter:1997:SWA

[563] Gail Letzter. Subalgebras which appear in quantum Iwasawa decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1206–1223, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Orsted:1997:TPA

[564] Bent Ørsted and Genkai Zhang. Tensor products of analytic continuations of holomorphic discrete series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1224–1241, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Randrianantoanina:1997:CSS

[565] Beata Randrianantoanina. 1-complemented subspaces of spaces with 1-unconditional bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1242–1264, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snaith:1997:HAC

[566] V. P. Snaith. Hecke algebras and class-group invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1265–1280, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sottile:1997:PFE

[567] Frank Sottile. Pieri’s formula via explicit rational equivalence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1281–1298, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tie:1997:ESN

[568] Jingzhi Tie. The explicit solution of the ¯∂-neumann problem in a non-isotropic Siegel domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):1299–1322, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sankaran:1997:SPP

[569] Parameswaran Sankaran and Peter Zvengrowski. Stable parallelizability of partially oriented flag manifolds II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49 (??):1323–1339, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:1997:AII

[570] Anonymous. Author index — index des auteurs — for 1997 — pour 1997. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??): 1340–1344, ???? 1997. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amberg:1998:SAG

[571] B. Amberg, O. Dickenschied, and Ya. P. Sysak. Subgroups of the adjoint group of a radical ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):3–15, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boroczky:1998:ASF

[572] Károly Böröczky, Jr. and Uwe Schnell. Asymptotic shape of finite packings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):16–28, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ding:1998:WNI

[573] Yong Ding and Shanzhen Lu. Weighted norm inequalities for fractional integral operators with rough kernel. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 29–39, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Englis:1998:GFP

[574] Miroslav Englis and Jaak Peetre.

Green’s functions for powers of the invariant Laplacian. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):40–73, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flicker:1998:EPF

[575] Yuval Z. Flicker. Elementary proof of the fundamental lemma for a unitary group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):74–98, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izuchi:1998:IST

[576] Keiji Izuchi and Yasuo Matsugu. A_{ϕ}- invariant subspaces on the torus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 99–133, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Medan:1998:CLS

[577] Christine Médan. On critical level sets of some two degrees of freedom integrable Hamiltonian systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):134–151, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Min:1998:IRF

[578] G. Min. Inequalities for rational functions with prescribed poles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):

152–166, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halverson:1998:MNR

[579] Tom Halverson and Arun Ram. Murnaghan–Nakayama rules for characters of Iwahori–Hecke algebras of the complex reflection groups G(r, p, n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):167–192, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xu:1998:IOH

[580] Yuan Xu. Intertwining operator and h-harmonics associated with reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):193–209, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhao:1998:IBG

[581] Kaiming Zhao. Isomorphisms between generalized Cartan type W Lie algebras in characteristic 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):210–224, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benkart:1998:DIF

[582] Georgia Benkart. Derivations and invariant forms of Lie algebras graded by finite root systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):225–241, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benoist:1998:IDS

[583] Joël Benoist. Intégration du sousdifférentiel proximal: un contre exemple. (French) [Integration of the proximal subdifferential: a counterexample]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):242–265, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Britten:1998:TFP

[584] D. J. Britten and F. W. Lemire. The torsion free Pieri formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):266–289, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davidson:1998:NDA

[585] Kenneth R. Davidson and Gelu Popescu.

Noncommutative disc algebras for semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):290–311, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokuchaev:1998:UGR

[586] Michael A. Dokuchaev and Maria Lucia Sobral Singer. Units in group rings of free products of prime cyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):312–322, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dykema:1998:PIS

[587] Kenneth J. Dykema and Mikael Rørdam. Purely infinite, simple C^{∗}- algebras arising from free product constructions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):323–341, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giraldo:1998:SFM

[588] Antonio Giraldo. Shape fibrations, multivalued maps and shape groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 342–355, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1998:SNU

[589] Leonard Gross. Some norms on universal enveloping algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):356–377, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kurth:1998:EPA

[590] Alexandre Kurth. Equivariant polynomial automorphism of Θ-representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):378–400, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:1998:HCU

[591] Yuanlin Li. The hypercentre and the n-centre of the unit group of an integral group ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):401–411, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McIntosh:1998:ATS

[592] Richard J. McIntosh. Asymptotic transformations of q-series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):412–425, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMullen:1998:GRS

[593] Peter McMullen. The groups of the regular star-polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):426–448, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aulaskari:1998:SRS

[594] Rauno Aulaskari, Yuzan He, Juha Ristioja, and Ruhan Zhao. Q_{p} spaces on Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):449–464, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balog:1998:SPA

[595] Antal Balog. Six primes and an almost prime in four linear equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 465–486, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barlow:1998:LPD

[596] Martin T. Barlow. On the Liouville property for divergence form operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):487–496, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bolle:1998:MIA

[597] Philippe Bolle. Morse index of approximating periodic solutions for the billiard problem. application to existence results. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):497–524, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brockman:1998:NOV

[598] William Brockman and Mark Haiman. Nilpotent orbit varieties and the atomic decomposition of the q-kostka polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):525–537, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Froese:1998:UBR

[599] Richard Froese. Upper bounds for the resonance counting function of Schrödinger operators in odd dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):538–546, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [699].

Gauthier:1998:MLT

[600] Paul M. Gauthier. Mittag-Leffler theorems on Riemann surfaces and Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):547–562, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldston:1998:PSS

[601] D. A. Goldston and C. Y. Yildirim. Primes in short segments of arithmetic progressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):563–580, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kamiyama:1998:HSP

[602] Yasuhiko Kamiyama. The homology of singular polygon spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):581–594, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luo:1998:MFC

[603] Donghan Luo and Thomas MacGregor.

Multipliers of fractional Cauchy transforms and smoothness conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 595–604, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guzman-Partida:1998:HSC

[604] Martha Guzmán-Partida and Salvador Pérez-Esteva. Hardy spaces of conjugate systems of temperatures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 605–619, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sjerve:1998:ETA

[605] Denis Sjerve and Qing Jie Yang. The Eichler trace of Z_{p} actions on Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):620–637, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strichartz:1998:FL

[606] Robert S. Strichartz. Fractals in the large. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):638–657, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Symesak:1998:HOP

[607] Frédéric Symesak. Hankel operators on pseudoconvex domains of finite type in C^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):658–672, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carey:1998:FMS

[608] Alan Carey and John Phillips. Fredholm modules and spectral flow. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 673–718, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gobel:1998:IAF

[609] Rüdiger Göbel and Saharon Shelah. Indecomposable almost free modules— the local case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):719–738, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Godsil:1998:EDR

[610] C. D. Godsil. Eigenpolytopes of distance regular graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):739–755, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brydges:1998:ERG

[611] D. Brydges, J. Dimock, and T. R.

Hurd. Estimates on renormalization group transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):756–793, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Louboutin:1998:UBA

[612] Stéphane Louboutin. Upper bounds on |L(1, χ)| and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 794–815, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mazorchuk:1998:TRG

[613] Volodymyr Mazorchuk. Tableaux realization of generalized Verma modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):816–828, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putcha:1998:CCN

[614] Mohan S. Putcha. Conjugacy classes and nilpotent variety of a reductive monoid. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):829–844, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scheerer:1998:LSC

[615] H. Scheerer and D. Tanré. Lusternik– Schnirelmann category and algebraic Rlocal homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):845–862, ????

1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yekutieli:1998:SFE

[616] Amnon Yekutieli. Smooth formal embeddings and the residue complex. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (4):863–896, August 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bloom:1998:FML

[617] Walter R. Bloom and Zengfu Xu. Fourier multipliers for local Hardy spaces on Chébli–Trimèche hypergroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):897–928, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Broer:1998:DVS

[618] Abraham Broer. Decomposition varieties in semisimple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 929–971, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruchert:1998:TCE

[619] Gerd Brüchert. Trace class elements and cross-sections in Kac–Moody groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):972–1006, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elder:1998:GMS

[620] G. Griffith Elder. Galois module structure of ambiguous ideals in biquadratic extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):1007–1047, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goerss:1998:LTS

[621] P. G. Goerss and J. F. Jardine. Localization theories for simplicial presheaves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):1048–1089, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lohoue:1998:TRG

[622] Noël Lohoué and Sami Mustapha. Sur les transformées de Riesz sur les groupes de Lie moyennables et sur certains espaces homogènes. (French) [On Riesz transforms on amenable Lie groups and on certain homogeneous spaces]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1090–1104, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1998:TRT

[623] Brooks Roberts. Tempered representations and the theta correspondence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):1105–1118, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anand:1998:WSI

[624] Christopher Kumar Anand. Ward’s solitons II: exact solutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):

1119–1137, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chalov:1998:CIM

[625] P. A. Chalov, T. Terzioglu, and V. P. Zahariuta. Compound invariants and mixed F -, DF-power spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1138–1162, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1998:GEH

[626] Jingyi Chen and Elton P. Hsu. Gradient estimates for harmonic functions on manifolds with Lipschitz metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1163–1175, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dobson:1998:IPM

[627] Edward Dobson. Isomorphism problem for metacirculant graphs of order a product of distinct primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):1176–1188, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Engler:1998:TRR

[628] Antonio José Engler. Totally real rigid elements and Galois theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1189–1208, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fukuma:1998:LBQ

[629] Yoshiaki Fukuma. A lower bound for K_{X}L of quasi-polarized surfaces (X, L) with non-negative Kodaira dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):1209–1235, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalton:1998:BLU

[630] N. J. Kalton and L. Tzafriri. The behaviour of Legendre and ultraspherical polynomials in L_{p}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1236–1252, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lopez-Bautista:1998:IRC

[631] Pedro Ricardo López-Bautista and Gabriel Daniel Villa-Salvador. Integral representation of p-class groups in Z_{p}-extensions and the Jacobian variety. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50 (??):1253–1272, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lubinsky:1998:MCL

[632] D. S. Lubinsky. Mean convergence of Lagrange interpolation for exponential weights on [−1, 1]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(6):1273– 1297, December 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Milson:1998:IGL

[633] Robert Milson. Imprimitively generated Lie-algebraic Hamiltonians and separation of variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):1298–1322, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morales:1998:LHW

[634] Jorge Morales. L’invariant de Hasse– Witt de la forme de Killing. (French) [The Hasse–Witt invariant in Killing form]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):1323–1336, ???? 1998. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:1998:AII

[635] Anonymous. Author index — index des auteurs — for 1998 — pour 1998. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??): 1337–1341, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allday:1999:CGK

[636] C. Allday and V. Puppe. On a conjecture of Goresky, Kottwitz and MacPherson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):3–9, February 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1999:TF

[637] M. Chacron, J.-P. Tignol, and A. R. Wadsworth. Tractable fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):

10–25, February 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabian:1999:SRS

[638] Marián Fabian and Boris S. Mordukhovich. Separable reduction and supporting properties of Fréchet-like normals in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1): 26–48, February 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ndombol:1999:AQC

[639] Bitjong Ndombol and M. El haouari.

Algèbres quasi-commutatives et carrés de Steenrod. (French) [Quasi-commutative algebras and Steenrod’s squares]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1): 49–68, February 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reichstein:1999:THJ

[640] Zinovy Reichstein. On a theorem of Hermite and Joubert. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):69–95, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosler:1999:PCS

[641] Margit Rösler and Michael Voit. Partial characters and signed quotient hypergroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):96–116, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sauer:1999:MFP

[642] A. Sauer. Meromorphic functions with prescribed asymptotic behaviour, zeros and poles and applications in complex approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):117–129, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Savin:1999:DPA

[643] Gordan Savin and Wee Teck Gan. The dual pair G_{2} × PU_{3}(D) (p-adic case). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (1):130–146, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Suarez:1999:HAM

[644] Daniel Suárez. Homeomorphic analytic maps into the maximal ideal space of H^{∞}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):147–163, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tan:1999:PSE

[645] Victor Tan. Poles of Siegel Eisenstein series on U(n, n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):164–175, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderPoorten:1999:VDE

[646] Alfred van der Poorten and Kenneth S. Williams. Values of the Dedekind eta function at quadratic irrationalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):176–224, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [698].

Betke:1999:AFL

[647] U. Betke and K. Böröczky, Jr. Asymptotic formulae for the lattice point enumerator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):225–249, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Combari:1999:CSC

[648] C. Combari, R. Poliquin, and L. Thibault.

Convergence of subdifferentials of convexly composite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):250–265, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deitmar:1999:SET

[649] Anton Deitmar and Werner Hoffman.

Spectral estimates for towers of noncompact quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):266–293, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Enochs:1999:HQM

[650] Edgar E. Enochs and Ivo Herzog. A homotopy of quiver morphisms with applications to representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):

294–308, April 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leung:1999:SSS

[651] Denny H. Leung and Wee-Kee Tang. Symmetric sequence subspaces of C(α), II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (2):309–325, April 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1999:ASO

[652] W. J. Martin and D. R. Stinson. Association schemes for ordered orthogonal arrays and (T, M, S)-nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):326–346, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muhlherr:1999:EMQ

[653] Bernhard Mühlherr and Hendrik Van Maldeghem. Exceptional Moufang quadrangles of type F_{4}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):347–371, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mytnik:1999:UCS

[654] Leonid Mytnik. Uniqueness for a competing species model. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):372–448, April 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bahn:1999:BMT

[655] Hyoungsick Bahn and Paul Ehrlich. A Brunn–Minkowski type theorem on the Minkowski spacetime. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):449–469, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bshouty:1999:EUH

[656] D. Bshouty and W. Hengartner. Exterior univalent harmonic mappings with finite Blaschke dilatations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3): 470–487, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burgess:1999:HAS

[657] W. D. Burgess and Manuel Saorín. Homological aspects of semigroup gradings on rings and algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):488–505, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elduque:1999:PIE

[658] A. Elduque and A. V. Iltyakov. On polynomial invariants of exceptional simple algebraic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):506–522, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabbri:1999:RVH

[659] Marc A. Fabbri and Frank Okoh. Representations of Virasoro–Heisenberg algebras and Virasoro-toroidal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (3):523–545, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felten:1999:SCI

[660] M. Felten. Strong converse inequalities for averages in weighted L^{p} spaces on [−1, 1]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):546–565, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferenczi:1999:QHI

[661] V. Ferenczi. Quotient hereditarily indecomposable Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3): 566–584, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mansfield:1999:SFD

[662] R. Mansfield, H. Movahedi-Lankarani, and R. Wells. Smooth finite dimensional embeddings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):585–615, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Panyushev:1999:PSA

[663] Dmitri I. Panyushev. Parabolic subgroups with Abelian unipotent radical as a testing site for invariant theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (3):616–635, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paul:1999:FOD

[664] Annegret Paul. First occurrence for the dual pairs (U(p, q), U(r, s)). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):

636–657, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shumyatsky:1999:NSL

[665] Pavel Shumyatsky. Nilpotency of some Lie algebras associated with p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (3):658–672, June 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barlow:1999:BMH

[666] Martin T. Barlow and Richard F. Bass.

Brownian motion and harmonic analysis on Sierpinski carpets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):673–744, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Echterhoff:1999:ICD

[667] Siegfried Echterhoff and John Quigg.

Induced coactions of discrete groups on C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):745–770, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flicker:1999:SBP

[668] Yuval Z. Flicker. Stable bi-period summation formula and transfer factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (4):771–791, August 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1999:TPT

[669] G. Grätzer and F. Wehrung. Tensor products and transferability of semilattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):792–815, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1999:NFS

[670] Brian C. Hall. A new form of the Segal– Bargmann transform for Lie groups of compact type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):816–834, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1999:LSM

[671] Henry H. Kim. Langlands–Shahidi method and poles of automorphic L-functions: Application to exterior square L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):835–849, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muhly:1999:TAI

[672] Paul S. Muhly and Baruch Solel. Tensor algebras, induced representations, and the Wold decomposition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4): 850–880, August 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Witherspoon:1999:RRC

[673] Sarah J. Witherspoon. The representation ring and the centre of a Hopf algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):881–896, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Astey:1999:CCP

[674] L. Astey, S. Gitler, E. Micha, and G. Pastor. Cohomology of complex projective Stiefel manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5): 897–914, October 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balogh:1999:QCP

[675] Zoltán M. Balogh and Christoph Leuenberger. Quasiconformal contactomorphisms and polynomial hulls with convex fibers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):915–935, October 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

David:1999:GRN

[676] Chantal David, Hershy Kisilevsky, and Francesco Pappalardi. Galois representations with non-surjective traces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (5):936–951, October 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deitmar:1999:LMS

[677] Anton Deitmar and Werner Hoffmann.

On limit multiplicities for spaces of automorphic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):952–976, October 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1999:EPN

[678] Stephen D. Fisher and Dmitry Khavinson. Extreme Pick–Nevanlinna interpolants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):977–995, October 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1999:CET

[679] Roger L. Jones, Joseph M. Rosenblatt, and Máté Wierdl. Counting in ergodic theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):996–1019, October 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kozlov:1999:FSM

[680] Dmitry N. Kozlov. On functions satisfying modular equations for infinitely many primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):1020–1034, October 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Litherland:1999:HAC

[681] R. A. Litherland. The homology of Abelian covers of knotted graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5): 1035–1072, October 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nielsen:1999:HPD

[682] Ole A. Nielsen. The Hausdorff and packing dimensions of some sets related to Sierpiński carpets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5): 1073–1088, October 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vakil:1999:CNQ

[683] Ravi Vakil. The characteristic numbers of quartic plane curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5): 1089–1120, October 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arnold:1999:FSL

[684] V. I. Arnold. First steps of local contact algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1123– 1134, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arthur:1999:EFC

[685] James Arthur. Endoscopic L-functions and a combinatorial identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1135– 1148, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1999:LGG

[686] A. M. Cohen, H. Cuypers, and H. Sterk.

Linear groups generated by reflection tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1149–1174, December 1999. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehrer:1999:RSU

[687] G. I. Lehrer and T. A. Springer. Reflection subquotients of unitary reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1175–1193, December 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lusztig:1999:SNE

[688] G. Lusztig. Subregular nilpotent elements and bases in K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1194– 1225, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKay:1999:SAC

[689] John McKay. Semi-affine Coxeter– Dynkin graphs and G ⊆ SU_{2}(C). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1226– 1229, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1999:STE

[690] Michael I. Hartley, Peter McMullen, and Egon Schulte. Symmetric tessellations on Euclidean space-forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1230– 1239, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Monson:1999:RRT

[691] B. Monson and A. Ivić Weiss. Realizations of regular toroidal maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1240– 1257, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baake:1999:SSR

[692] Michael Baake and Robert V. Moody.

Similarity submodules and root systems in four dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1258– 1276, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1999:IIP

[693] G. C. Shephard. Isomorphism invariants for projective configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1277– 1299, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conway:1999:ESS

[694] J. H. Conway, E. M. Rains, and N. J. A. Sloane. On the existence of similar sublattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1300– 1306, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1999:QIC

[695] Norman W. Johnson and Asia Ivić Weiss. Quadratic integers and Coxeter groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1307–1336, December 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:1999:AII

[696] Anonymous. Author index — index des auteurs — for 1999 — pour 1999. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1337– 1341, December 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edward:2000:CST

[697] Julian Edward. Corrigendum to “Spectral Theory for the Neumann Laplacian on Planar Domains with Horn-Like Ends”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1):119–122, February 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [516].

vanderPoorten:2001:VDE

[698] Alfred J. van der Poorten and Kenneth S. Williams. Values of the Dedekind eta function at quadratic irrationalities: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 434–448, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [646].

Froese:2001:CUB

[699] Richard Froese. Correction to: “Upper Bounds for the Resonance Counting Function of Schrödinger Operators in Odd Dimensions”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):756–757, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [599].

Stahl:2008:EZS

[700] Saul Stahl. Erratum: “On the Zeros of Some Genus Polynomials”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):960, ????

2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See